### v1

# On the complexity of estimating ground state entanglement and free energy

Abstract
Understanding the entanglement structure of local Hamiltonian ground spaces is a physically motivated problem, with applications ranging from tensor network design to quantum error-correcting codes. To this end, we study the complexity of estimating ground state entanglement, and more generally entropy estimation for low energy states and Gibbs states. We find, in particular, that the classes $\mathsf{qq\mhyphen QAM}$ [Kobayashi, le Gall, Nishimura, SICOMP 2019] (a quantum analogue of public-coin AM) and $\mathsf{QMA}(2)$ (QMA with unentangled proofs) play a crucial role for such problems, showing: (1) Detecting a high-entanglement ground state is $\mathsf{qq\mhyphen QAM}$-complete, (2) computing an additive error approximation to the Helmholtz free energy (equivalently, a multiplicative error approximation to the partition function) is in $\mathsf{qq\mhyphen QAM}$, (3) detecting a low-entanglement ground state is $\mathsf{QMA}(2)$-hard, and (4) detecting low energy states which are close to product states can range from QMA-complete to $\mathsf{QMA}(2)$-complete. Our results make progress on an open question of [Bravyi, Chowdhury, Gosset and Wocjan, Nature Physics 2022] on free energy, and yield the first $\mathsf{QMA}(2)$-complete Hamiltonian problem using local Hamiltonians (cf. the sparse $\mathsf{QMA}(2)$-complete Hamiltonian problem of [Chailloux, Sattath, CCC 2012]).
1 Introduction
Estimating low energy properties of $k$-local Hamiltonian systems is a central challenge of quantum many-body physics. Here, a $k$-local Hamiltonian is an $n$-qubit Hermitian operator $H$ specified succinctly via $H=\sum_{i}H_{i}$, where each $H_{i}$ acts non-trivially on only $O(1)$ qubits. Such operators generalize classical Boolean $k$-local formulae, such as in $k$-SAT, and like their classical Boolean counterparts, $k$-local Hamiltonians are a cornerstone of quantum complexity theory [Osb12, GHLS15]. Briefly, low energy properties of $k$-local Hamiltonians are analogous to properties of the solution space to $k$-local Boolean formulae.
The “canonical” low-energy property of $k$-local Hamiltonians is their ground state energy, whose estimation is complete for Quantum Merlin Arthur [^2] QMA is the canonical quantum generalization of NP [AN02, Gha24]. (QMA) [KSV02, KR03, KKR06], and which generalizes the maximum number of satisfiable clauses of a Boolean formula. Formally, the ground state energy of $H$ is its smallest eigenvalue $\lambda_{\min}(H)$, which corresponds to the energy level the system relaxes into when cooled to low temperature; this quantity plays an important role, e.g., in quantum chemistry in estimating reaction rates. However, what is also often important (e.g. for quantum error-correcting codes) are properties of the corresponding ground space itself, i.e. eigenvectors/ground states $\left|\psi\right\rangle$ such that $H\left|\psi\right\rangle=\lambda_{\min}(H)\left|\psi\right\rangle$. Ground states encode the state the system can take on at its ground energy.
To this end, numerous physically motivated low energy properties beyond ground state energy have been complexity theoretically characterized over the years, including ground state degeneracy [BFS11, SZ09], simulation of local observables [Amb14, GY18, GPY20, GR22, WBG23], spectral gaps [Amb14, CPW15, GY18], energy barriers [GS18, GMV17, GR23, GH24], phase transitions [WB21, PLC24], and “universal” Hamiltonians whose low energy space can reproduce the physics of other many-body systems [BH17, CMP18, KPBC22]. Until recently, however, surprisingly absent from this list was the question:
What is the complexity of deciding whether $H$ has an entangled ground state?
In this work, we will show that this question is more nuanced than one might expect, and draw connections to various complexity classes and related important concepts such as the free energy of Gibbs states.
Ground state entanglement: Prior work.
We begin by discussing three recent works which have studied the complexity of detecting ground state entanglement. The first two [GH24a, Bou+24] achieve hardness for the Learning With Errors (LWE) problem [Reg09], and the third [GH24] for Quantum Classical Merlin Arthur (QCMA), i.e. QMA but with a classical proof. In more detail, Gheorghiu and Hoban [GH24a] first showed that given two local Hamiltonians and a pre-specified cut $A$ versus $B$ among qubits, it is LWE-hard to decide: Among all ground states $\left|\psi_{1}\right\rangle$ of $H_{1}$ and $\left|\psi_{2}\right\rangle$ of $H_{2}$ of minimal entanglement entropy, which of $\left|\psi_{1}\right\rangle$ or $\left|\psi_{2}\right\rangle$ has larger entanglement entropy [^3] Entanglement entropy is a standard entanglement measure, defined for a bipartite pure state $\left|\psi\right\rangle_{AB}$ as $S(\rho_{A})$, for reduced state $\rho_{A}:=\operatorname{Tr}_{B}(\rho_{AB})$ and $S(\rho):=-\operatorname{Tr}(\rho\log\rho$) the von Neumann entropy. across $A$ versus $B$? Next, Bouland, Fefferman, Ghosh, Metger, Vazirani, Zhang and Zhou [Bou+24, BZZ24] showed that deciding whether the ground states of geometrically constrained $H$ have volume law or area law entanglement across a given set of cuts is LWE-hard. Note that the input model in [Bou+24] is cryptographic in flavor, in that $H$ is drawn randomly from one of two computationally indistinguishable families. Finally, in the standard complexity theoretic input model of a single $H$ being given as input, Gharibian and Hecht [GH24] showed: Given $H$, cut $A$ versus $B$, and thresholds $\eta_{1}>\eta_{2}$, it is QCMA-hard to distinguish whether $H$ has a ground state with entanglement entropy at most $\eta_{1}$ across the $A$ versus $B$ cut, or whether all ground states have entanglement entropy at least $\eta_{2}$ across this cut, even if $\eta_{2}/\eta_{1}$ is large. Thus, entanglement entropy estimation is even QCMA-hard to approximate.
Our starting point: Entropy estimation.
A challenge in studying ground state entanglement is that given a physical copy of a bipartite state $\left|\psi\right\rangle_{AB}$, estimating entropy $S(\psi_{A})$ is not known to be efficiently computable quantumly. This is why the careful reader may notice that, e.g., [GH24] does not show containment in any complexity class, but only QCMA-hardness. Indeed, it is generally believed entropy estimation is not in BQP. If $\left|\psi\right\rangle_{AB}$ were specified via a poly-size quantum preparation circuit and given threshold $t$, deciding whether $S(\psi_{A})$ is at least $t+1/2$ or at most t1/2t-1/2 is complete for Non-Interactive Quantum Statistical Zero Knowledge ($\mathsf{NIQSZK}$) [Kob03, BST08]! [^4] This result is sort of implicit in the literature. In [BST08] it is proven that estimating entropy is contained in $\mathsf{QSZK}$. Closer inspection reveals that they reduce to determining the distance between some state $\sigma$, prepared by a poly-size circuit, and the maximally mixed state $\widetilde{I}$, which is proven to be $\mathsf{NIQSZK}$-complete in [Kob03]. Note that while $\mathsf{NIQSZK}\subseteq\mathsf{QSZK}\subseteq\mathsf{QIP}(2)\subseteq\mathsf{PSPACE}$, for $\mathsf{QIP}(2)$ the class of decision problems with $2$-round [^5] While $\mathsf{QIP}(3)=\mathsf{QIP}=\mathsf{PSPACE}$ [KW00, JJUW11], $\mathsf{QIP}(2)$ remains a poorly understood class. quantum interactive proofs, the relationship between $\mathsf{QMA}$ and $\mathsf{NIQSZK}$ is unclear. In fact, there is an oracle separation between $\mathsf{NIQSZK}$ and $\mathsf{PP}$[BCHTV19].
In the case of ground state entanglement, we face the additional challenge that we do not even have a poly-size circuit preparing $\left|\psi\right\rangle_{AB}$ (unless $\mathsf{QMA}=\mathsf{QCMA}$) — this breaks even the QSZK containment proof of [BST08] for entropy estimation, raising various questions: What is the best upper bound on ground state entanglement estimation? Does the complexity change if we are interested in high or low entanglement ground states? And is there a connection to $\mathsf{QMA}(2)$, which is $\mathsf{QMA}$ with unentangled proofs?
Beyond the low energy setting: Free energy of Gibbs states.
Finally, our study will also take us beyond the low energy setting to high temperature Gibbs states, where entropy estimation again plays an important role. Specifically, a Gibbs state for a given Hamiltonian $H$ is the mixed state $\rho_{\beta}=e^{-\beta H}/\mathcal{Z}$, for normalization $\mathcal{Z}:=\operatorname{Tr}(e^{-\beta H})$ denoting the quantum partition function. Roughly, at temperature $T=1/\beta$, the quantum system thermalizes to $\rho_{\beta}$, where the ground state is recovered by setting $T=0$. The relevant quantity for us is the Helmholtz free energy, which plays an important role (e.g.) in mapping out phase diagrams, and which generalizes ground state energy to temperatures $T>0$. Specifically, the free energy of $H$ at inverse temperature $\beta$ is $\mathcal{F}=\operatorname{Tr}(\rho_{\beta}H)-(1/\beta)S(\rho_{\beta})$, where note the entropy $S(\rho_{\beta})$.
Bravyi, Chowdhury, Gosset and Wocjan [BCGW22] studied the natural question of additive error approximations to $\mathcal{F}$ using an equivalent definition,
$\mathcal{F}=-(1/\beta)\log\mathcal{Z}.$ (1)
Thus, an additive approximation of $\mathcal{F}$ is equivalent to a relative approximation of $\mathcal{Z}$, dubbed the Quantum Partition Function (QPF) problem in [BCGW22]. Classically, although exactly computing the partition function is $\textup{P}-$hard, relative approximation is in $\mathsf{BPP}^{\mathsf{NP}}$ via Stockmeyer counting. In contrast, in the quantum case, it is only known that that the problem is contained in $\mathsf{P}^{\#\mathsf{P}}$ [BCGW22, BFS11], at least QMA-hard due to the special case $T=0$, and if $\beta$ is complex, then QPF is $\textup{P}-$hard within constant relative error (in fact, even for classical $H$ [GG17, FM17]). In this direction, [BCGW22] showed that QPF is equivalent under poly-time reductions to various problems, such as the number of eigenvalues of a $k$-local Hamiltonian in a given energy interval, but the complexity of these equivalent problems is also unfortunately unknown. This brings us to an open question of [BCGW22], which we also focus on here: What is the complexity of QPF, i.e. additive approximations to the free energy?
Our results.
We study three computational problems involving low energy entanglement, as well as free energy estimation (QPF). We group our results by the relevant complexity class.
1. High entanglement low energy states and free energy: $\mathsf{qq\mhyphen QAM}$.
[…]
1.1 High entanglement low energy states. With $\mathsf{qq\mhyphen QAM}$ in hand, the first problem we study is:
Definition 1.1 (High Entropy Low Energy State ($\mathsf{HELES}$)).
Given is a $k$-local Hamiltonian $H_{AB}=\sum_{j=1}^{m}H_{j}$ on registers $A$ and $B$ of $n_{A}+n_{B}=n$ qubits, energy threshold parameters $\alpha,\beta$ such that $\beta-\alpha=1/\operatorname{poly}(n)$ and entropy threshold parameters $s,t$ such that $s-t=1/\operatorname{poly}(n)$. Decide:
  • (YES case) If $\exists\;\left|\psi\right\rangle_{AB}$ such that $\left\langle\psi\right|H\left|\psi\right\rangle\leq\alpha$ and $S(\psi_{A})\geq s$.
  • (NO case) If $\forall\;\left|\psi\right\rangle_{AB}$ either $\left\langle\psi\right|H\left|\psi\right\rangle\geq\beta$ or $S(\psi_{A})\leq t$.
Above and in this work, we use shorthand $\psi_{A}$ to denote $\operatorname{Tr}_{B}(\lvert\psi\rangle\langle\psi\rvert_{AB})$. In words, HELES asks: Does $H$ have a high entanglement low energy state across the $A$ versus $B$ cut? (We warn the reader already that whether the YES case corresponds to a high or low entanglement state will change the complexity analysis completely, to be discussed shortly.) Our first result is as follows.
Theorem 1.2.
$\mathsf{HELES}$ is $\mathsf{qq\mhyphen QAM}$-complete for $k\geq 5$. Hardness remains even when restricting to physically motivated Hamiltonians such as the 2D Heisenberg model (Corollary 8.6).
Thus, while entanglement estimation is NIQSZK-complete when $\left|\psi\right\rangle_{AB}$ is specified via quantum circuit [Kob03, BST08], the complexity jumps to $\mathsf{qq\mhyphen QAM}$ when $\left|\psi\right\rangle_{AB}$ is specified implicitly via a local Hamiltonian’s low energy space. We say “jumps” because $\mathsf{NIQSZK}\subseteq\mathsf{qq\mhyphen QAM}$.
1.2 Free energy. We next study the complexity of additive error approximations to free energy (equivalent to QPF of [BCGW22]).
Definition 1.3.
(Free Energy Approximation ($\mathsf{FEA}$)) Given is a $k$-local Hamiltonian $H$ on $n$ qubits, inverse temperature $\beta$, and two parameters $a,b$ satisfying $b-a=1/\operatorname{poly}(n)$. Decide:
  • (YES case) The free energy of $H$ at inverse temperature $\beta$ is less than $a$. That is $\mathcal{F}(H)\leq a$.
  • (NO case) The free energy of $H$ at inverse temperature $\beta$ is larger than $b$. That is, $\mathcal{F}(H)\geq b$.
We show:
Theorem 1.4.
for all $O(1)$-local Hamiltonians, $\mathsf{FEA}$ is in $\mathsf{qq\mhyphen QAM}$.
Thus, we obtain a bound on the power of additive approximations to the free energy, or equivalently, relative error approximations to a quantum partition function. It is unclear however, how $\mathsf{qq\mhyphen QAM}$ relates to the best previously known bound $\mathsf{P}^{\#\mathsf{P}}$.
2. Low entanglement and approximately product ground states: $\mathsf{QMA}(2)$.
We tie the complexity of our next two problems to $\mathsf{QMA}(2)$ [KMY03], a class notorious for resisting anything but non-trivial complexity bounds $\mathsf{QMA}\subseteq\mathsf{QMA}(2)\subseteq\mathsf{NEXP}$.
2.1 Low entanglement low energy states. Our third problem is a seemly innocuous variant of HELES, in which we seek a low entanglement state (as opposed to a high entanglement one).
Definition 1.5.
(Low Entropy Low Energy State ($\mathsf{LELES}$)) Given is a $k$-local Hamiltonian $H$ on two registers $A$ and $B$ of $n_{A}+n_{B}=n$ qubits, energy thresholds $\alpha,\beta$ such that $\beta-\alpha=1/\operatorname{poly}(n)$, and entropy thresholds $s,t$ such that $s-t\geq\frac{1}{\operatorname{poly}(n)}$. Decide:
  • (YES case) If $\exists\;\left|\psi\right\rangle_{AB}$ such that $\left\langle\psi\right|H\left|\psi\right\rangle\leq\alpha$ and $S(\psi_{A})\leq t$.
  • (NO case) If $\forall\;\left|\psi\right\rangle_{AB}$ either $\left\langle\psi\right|H\left|\psi\right\rangle\geq\beta$ or $S(\rho_{A})\geq s$.
Note that $\mathsf{LELES}$ is not the complement of $\mathsf{HELES}$, as we have only flipped the condition on entropy. Formally, LELES is equivalent to the ground state entanglement problem shown QCMA-hard to approximate in [GH24]. Our initial aim was to show a non-trivial upper bound on LELES, which proved difficult; the following theorem explains why.
Theorem 1.6.
For $k\geq 5$, $\mathsf{LELES}$ is $\mathsf{QMA}(2)$-hard, even when $s-t=\Omega(n)$. Hardness remains even when restricting to physically motivated Hamiltonians such as the 2D Heisenberg model (Corollary 8.6).
Thus, upper bounding the complexity of LELES is at least as difficult as bounding the power of $\mathsf{QMA}(2)$, a longstanding open problem. Whether this implies finding low entanglement ground states is easier (LELES) or harder than finding high entanglement ground states (HELES) depends on what one believes regarding the power of $\mathsf{QMA}$ versus $\mathsf{QMA}(2)$. It is entirely plausible that $\mathsf{QMA}(2)=\mathsf{QMA}\subsetneq\mathsf{qq\mhyphen QAM}$ (i.e. LELES is easier than HELES), or that $\mathsf{QMA}(2)=\mathsf{NEXP}\supsetneq\mathsf{qq\mhyphen QAM}$ (i.e. HELES is much harder than LELES).
2.2 Low energy approximate product states. As entropy is a challenging quantity, and is a barrier to showing $\mathsf{QMA}(2)$ containment of $\mathsf{LELES}$, our fourth and final problem studies a different measure of entanglement: the trace distance to a product state.
Definition 1.7.
(Low Energy Approximate Product State ($\mathsf{LEAPS}$)) Given is a $k$-local Hamiltonian $H$ on $n$ qubits, energy thresholds $\alpha,\beta$, and distance thresholds $a,b<1$. Decide:
  • (YES case) There exists a low-energy state that is $a$-close to product in trace distance: $\exists\left|\psi\right\rangle_{AB},\left|\phi_{L}\right\rangle_{A},\left|\phi_{R}\right\rangle_{B}$ such that $\left\langle\psi\right|H\left|\psi\right\rangle\leq\alpha$ and $\|\lvert\psi\rangle\langle\psi\rvert_{AB}-\lvert\phi_{L}\rangle\langle\phi_{L}\rvert_{A}\otimes\lvert\phi_{R}\rangle\langle\phi_{R}\rvert_{B}\|_{1}\leq a$.
  • (NO case) All low-energy states are $b$-far from product: $\forall\left|\psi\right\rangle_{AB},\left|\phi_{L}\right\rangle_{A},\left|\phi_{R}\right\rangle_{B}$, either $\left\langle\psi\right|H\left|\psi\right\rangle\geq\beta$ or $\|\lvert\psi\rangle\langle\psi\rvert_{AB}-\lvert\phi_{L}\rangle\langle\phi_{L}\rvert_{A}\otimes\lvert\phi_{R}\rangle\langle\phi_{R}\rvert_{B}\|_{1}\geq b$.
[…]
Theorem 1.8.
$\mathsf{LEAPS}$ is $\mathsf{QMA}$-complete if $\beta-\alpha-a\|H\|_{\infty}\geq\frac{1}{\operatorname{poly}(n)}$ and $b>a$.
Theorem 1.9.
For all constant $k\geq 5$, $\mathsf{LEAPS}$ is $\mathsf{QMA}(2)$-complete if $\beta-\alpha=\frac{1}{\operatorname{poly}(n)}$, $b-a=\frac{1}{\operatorname{poly}(n)}$, $a=\frac{1}{\operatorname{poly}(n)}$ and $\beta=O(a^{6})$. Hardness remains even when restricting to physically motivated Hamiltonians such as the 2D Heisenberg model (Corollary 8.6).
Note that $\|H\|_{\infty}\geq 1$. Therefore, our $\mathsf{QMA}$-containment result requires $\beta>a$, whereas we only show $\mathsf{QMA}(2)$-hardness when $\beta=O(a^{6})$. In particular, as $a<1$, there is no overlap between the regimes where we show $\mathsf{QMA}$ and $\mathsf{QMA}(2)$-completeness.
Thus, we obtain a natural local Hamiltonian problem whose complexity “leaps” from $\mathsf{QMA}$ to $\mathsf{QMA}(2)$, depending on the choice of parameters. LEAPS is additionally appealing for our context as it studies not just exact product states (as in [CS12]), but the more robust notion of close to product states, which corresponds to low entropy.
Techniques.
We make use of two main technical tools, which we encapsulate into two standalone lemmas (Lemma 3.1 and Lemma 3.2, respectively) for ease of further adoption. The first is a $\mathsf{qq\mhyphen QAM}$ protocol for entropy verification, which we use to prove that $\mathsf{qq\mhyphen QAM}$ contains $\mathsf{HELES}$ and $\mathsf{FEA}$. This protocol is not new; it appears implicitly in [KGN19, proof of Lemma 33] and is based on ideas from [BST08, Section 5]. The second tool is a channel-to-Hamiltonian construction, which we use to prove that $\mathsf{HELES}$ is $\mathsf{qq\mhyphen QAM}$-hard and to prove that $\mathsf{LELES}$ and $\mathsf{LEAPS}$ are $\mathsf{QMA}(2)$-hard. We now sketch each of these tools.
Entropy verification. The workhorse of this protocol is a quantum extractor, i.e. a channel $T$ roughly satisfying $T(\rho)\approx\widetilde{I}$ if and only if the min-entropy [^8] The min-entropy is defined as $H_{\infty}(\rho)=-\log\|\rho\|_{\infty}$ and lower bounds the usual Von Neumann entropy. $H_{\infty}(\rho)$ is sufficiently large (by $\widetilde{I}$ we denote the maximally mixed state). We will use the Stinespring dilation unitary $U_{T}$ of $T$, which satisfies $T(\rho)=\operatorname{Tr}_{E}\left(U_{T}(\rho\otimes\lvert 0^{2n_{A}}\rangle\langle 0^{2n_{A}}\rvert_{E})U_{T}^{\dagger}\right)$. To help understand the protocol, one can adopt the view of $\mathsf{qq\mhyphen QAM}$ as QMA with a normalized Choi matrix $\rho_{AB}$ as proof. Recall this gave us the key property that $\rho_{A}=\widetilde{I}_{A}$. Note that by unitary freedom of purifications, the prover can prepare any desired purification of the $\widetilde{I}_{A}$ system.
To now check entropy, the verifier proceeds as follows (Figure 1). First, they “undo” $T$. That is, they apply $U_{T}^{\dagger}$ to the entire proof state $\rho_{AB}$. Next, they measure $B$ in the computational basis and reject if the result is not all 0’s. If this test passes with high probability, then the postmeasurement state $\sigma_{A}$ on register $A$ has high min-entropy. The reason for this is as follows. Suppose the probability of getting all 0’s was high. Then, the state after applying $U_{T}^{\dagger}$ is approximately $\sigma_{A}\otimes\lvert 0^{2n_{A}}\rangle\langle 0^{2n_{A}}\rvert_{B}$. Note that $U_{T}(\sigma_{A}\otimes\lvert 0^{2n_{A}}\rangle\langle 0^{2n_{A}}\rvert_{B})U_{T}^{\dagger}\approx\rho_{AB}$. In particular, if one would “redo” $T$, one gets $T(\sigma)=\operatorname{Tr}_{B}(U_{T}(\sigma_{A}\otimes\lvert 0^{2n_{A}}\rangle\langle 0^{2n_{A}}\rvert_{B})U_{T}^{\dagger})\approx\widetilde{I}_{A}$. Since $T$ is an extractor, this implies the min-entropy of $\sigma_{A}$ is large.
There is one more hurdle to overcome: we need to move from min-entropy to Von Neumann entropy. Note that the min-entropy lower bounds the Von Neumann entropy, so if the protocol accepts we can be sure that $\sigma_{A}$ has high Von Neumann entropy as well. What goes wrong is that if $S(\sigma_{A})$ is high, that does not imply that $T(\sigma)\approx\widetilde{I}$. Therefore, there are states that do have high Von Neumann entropy, but cannot be the output of the protocol (when the measurement accepts with high probability). To fix this, we use a flattening lemma, which states that $H_{\infty}(\rho^{\otimes q})\approx qS(\rho)$. It follows that is $S(\sigma)$ is high, then $\sigma^{\otimes q}$ can be the output of the protocol. Vice-versa, if $\sigma_{A^{\otimes q}}$ is the output of the protocol, the verifier knows that $S(\sigma_{A^{\otimes q}})$ is high. While the different copies of the $A$ register could be horribly entangled, subadditivity of Von Neumann entropy ensures that most of the reduced state $\sigma_{A_{i}}$ must also have high entropy. Curiously, this is where the potentially large complexity difference between $\mathsf{HELES}$ and $\mathsf{LELES}$ could come from: by subadditivity, the prover cannot entangle low entropy reduced states $\sigma_{A_{i}}$ to create a high entropy global state $\sigma_{A^{\otimes q}}$, but they can entangle high entropy reduced states to create a low entropy global state. This allows the prover to cheat in the NO case if one tries to run the entropy verification protocol to decide $\mathsf{LELES}$ in $\mathsf{qq\mhyphen QAM}$.
   $\rho_{A}=\widetilde{I}_{A}$ $U_{T}^{\dagger}$ $U_{T}^{\dagger}$ $\sigma_{A}$ $\rho_{B}$ $\approx\left|0^{2n_{A}}\right\rangle_{E}$
$\iff$    $\sigma_{A}$ $U_{T}$ $U_{T}$ $T(\sigma)\approx\widetilde{I}_{A}$ $\left|0^{2n_{A}}\right\rangle_{E}$ $\approx\rho_{B}$ (Traced out) $\iff H_{\infty}(\sigma)\geq k$
Figure 1: Schematic representation of the entropy verification protocol. If the proof is such that the $E$ register after applying $U_{T}^{\dagger}$ is approximately $|0^{2n_{A}}\rangle$, then the contents of the $A$ register $\sigma_{A}$ are such that $T(\sigma_{A})\approx\widetilde{I}$. As $T$ is an extractor, this implies that the min-entropy of $\sigma_{A}$ is large.
Channel-to-Hamiltonian construction. We next discuss our channel-to-Hamiltonian construction, which allows us to reduce hardness results for channel problems to hardness results for Hamiltonians. Specifically, we use it to prove that $\mathsf{HELES}$ is $\mathsf{qq\mhyphen QAM}$-hard and to that $\mathsf{LELES}$ and $\mathsf{LEAPS}$ are $\mathsf{QMA}(2)$-hard. This construction bypasses the “sparse Hamiltonian roadblock” of [CS12], allowing us to obtain $\mathsf{QMA}(2)$-hardness for $\mathsf{LELES}$ and $\mathsf{LEAPS}$ even for local $H$. This is because it allows us to reduce from the channel-based $\mathsf{QMA}(2)$-complete Pure Product Isometry Output (PPIO) problem of Gutoski, Hayden, Milner and Wilde [GHMW13], rather than utilizing the Product State Test of Harrow and Montanaro [HM13] as done in [CS12], which is a non-local operation.
The idea behind the channel-to-Hamiltonian construction is as follows. Given a channel $\Phi$, we wish to define a Hamiltonian whose ground space consists of all states the channel could output. To achieve this, we use Kitaev’s circuit-to-Hamiltonian construction [KSV02] applied to the Stinespring dilation unitary $U_{\Phi}$, which recall satisfies $\Phi(\rho)=\operatorname{Tr}_{E}(U_{\Phi}(\rho\otimes\lvert 0^{2n_{A}}\rangle\langle 0^{2n_{A}}\rvert_{E})U_{\Phi}^{\dagger})$. We make two modifications: we do not use an output penalty and we post idle the computation.
As we do not use an output penalty, the zero-energy space of the resulting Hamiltonian $H_{\Phi}$ will be the span of history states of the form
$\left|\psi_{\textup{hist}}\right\rangle$ $=\frac{1}{\sqrt{T+1}}\sum_{t=0}^{T}V_{t}\dots V_{1}\left|\psi\right\rangle_{A}\left|0^{n_{B}}\right\rangle_{B}\left|t\right\rangle_{C},$ (2)
[…]
$\left|\psi_{\textup{hist}}\right\rangle=\frac{1}{\sqrt{T+L+1}}\sum_{t=0}^{T-1}V_{t}\dots V_{1}\left|\psi\right\rangle_{A}\left|0^{n}\right\rangle_{B}\left|t\right\rangle_{C}+U_{\Phi}\left|\psi\right\rangle_{A}\left|0^{n}\right\rangle_{B}\left(\frac{1}{\sqrt{T+L+1}}\sum_{t=T}^{T+L}\left|t\right\rangle_{C}\right).$ (3)
[…]
“Leaping” from $\mathsf{QMA}$ to $\mathsf{QMA}(2)$. The crucial insight that allows us to prove that $\mathsf{LEAPS}$ “leaps” from $\mathsf{QMA}$-completeness to $\mathsf{QMA}(2)$-completeness is as follows. Consider a state $\left|\psi\right\rangle_{AB}$ that is $a$ close to $\left|\phi_{L}\right\rangle\left|\phi_{R}\right\rangle$ and has low energy. Suppose the $a$ parameter is substantially smaller than the gap $\beta-\alpha$. Then, the energy of $\left|\phi_{L}\right\rangle\left|\phi_{R}\right\rangle$ will not be too high, specifically it will be smaller than (say) $\alpha+\frac{\beta-\alpha}{4}$. To prove that $\left|\psi\right\rangle_{AB}$ is low energy and close to product, the prover can thus send the state $\left|\phi_{L}\right\rangle\left|\phi_{R}\right\rangle$, which the techniques from [CS12] allow them to do. The verifier can then check that $\left|\phi_{L}\right\rangle\left|\phi_{R}\right\rangle$ has energy $\leq\alpha+\frac{\beta-\alpha}{4}$ which will convince them that they are dealing with a $\mathsf{LEAPS}$ YES case. Note that in the No case, in particular all product states have energy at least $\beta$, so the prover cannot cheat. In other words, if $a$ is sufficiently smaller than $\beta$, the energy and distance checks can be combined into one.
Suppose now that $\beta$ is lot smaller than $a$. In this case it is entirely possible that $\left|\psi\right\rangle_{AB}$ has low energy and is $a$ close to product, but that any product state it is close to has energy $\geq\beta$. The energy and distance checks can thus not be combined like before. Instead, they need to be checked separately. This can indeed be done in $\mathsf{QMA}(2)$: the provers can send $\left|\psi\right\rangle_{AB}$ and the product state $\left|\phi_{L}\right\rangle\left|\phi_{R}\right\rangle$ close to it. As the verifier is promised that these are all in product, they can check the distance with a SWAP test. Intuitively, the reason for $\mathsf{QMA}(2)$ hardness is that the verifier needs to verify this product structure.
Open questions.
This work has singled out two classes, $\mathsf{qq\mhyphen QAM}$ and $\mathsf{QMA}(2)$, as being important to the study of entanglement of low energy states and free energy of Gibbs states. Some problems we have been able to characterize completely, such as $\mathsf{qq\mhyphen QAM}$-completeness of high energy ground state detection (Theorem 1.2) and $\mathsf{QMA}(2)$-completeness of low energy approximate product states verification (Theorem 1.9). On the other hand, the three main open questions we leave are: (1) What is a tight upper bound on the complexity of low energy ground state detection (LELES)? On the one hand, showing a bound such as PSPACE would entail a major breakthrough in quantum complexity theory, due to our $\mathsf{QMA}(2)$-hardness result (Theorem 1.6). On the other hand, a “lower” bar might be to show containment in $\mathsf{QMA}(2)$ itself, though it is unclear how one would verify entropy in $\mathsf{QMA}(2)$. (2) Can one show a lower bound of $\mathsf{qq\mhyphen QAM}$ on the complexity of additive error free energy/relative error partition function approximation, matching our containment result in $\mathsf{qq\mhyphen QAM}$ (Theorem 1.4)? A potential obstacle here is that we do not have a particularly clean method of reducing hard problems to Gibbs states (as opposed to low energy states, for which we can use Kitaev’s circuit-to-Hamiltonian construction). (3) What is the complexity of $\mathsf{LEAPS}$ between where we show $\mathsf{QMA}$-completeness and where we show $\mathsf{QMA}(2)$-completeness? That is, is there a “sharp” transition between $\mathsf{QMA}$ and $\mathsf{QMA}(2)$ or are there values where the problem is complete for some class between $\mathsf{QMA}$ and $\mathsf{QMA}(2)$?
Organization.
We begin with preliminaries in Section 2. Section 3 fleshes out our main two technical tools. Section 4 studies HELES, Section 5 for FEA, Section 6 for LELES, and Section 7 for LEAPS. Finally, we consider physical Hamiltonian in Section 8.
2 Preliminaries
Notation.
We use $\|\left|\psi\right\rangle\|_{2}$ to denote the Euclidean norm of $\left|\psi\right\rangle$, $\|\rho\|_{1}=\operatorname{Tr}(\sqrt{\rho\rho^{\dagger}})$ the trace norm of $\rho$ and $\|A\|_{\infty}=\max_{\left|\psi\right\rangle}\|A\left|\psi\right\rangle\|_{2}/\|\left|\psi\right\rangle\|_{2}$ the spectral/operator norm of $A$.
Throughout this paper, $\widetilde{I}\coloneq\frac{1}{d}I$ refers to the maximally mixed state on a Hilbert space of dimension $d$. The corresponding Hilbert space will either be clear from context, or denoted in a subscript, e.g., $\widetilde{I}_{A}$ for the maximally mixed state on $\mathcal{H}_{A}$.
When $\left|\psi\right\rangle$ is a pure state, we will use the shorthand notation $\psi\coloneq\lvert\psi\rangle\langle\psi\rvert$ for the corresponding density matrix. Note that the symbols $\psi,\phi$ are used for pure states only.
2.1 Complexity classes and problems
The class $\mathsf{qq\mhyphen QAM}$.
One of the main complexity classes used in this work is the relatively obscure $\mathsf{qq\mhyphen QAM}$, first defined in [KGN19]. This class is a quantum generalization of the classical $\mathsf{AM}$ but is potentially different from $\mathsf{QAM}$, the more common quantum generalization of $\mathsf{AM}$. We now sketch the difference between $\mathsf{qq\mhyphen QAM},\mathsf{QAM}$ and $\mathsf{AM}$.
$\mathsf{AM}$ can be defined as the class of languages accepted by protocols of the following form: First, a $\mathsf{BPP}$ verifier Arthur and a prover Merlin share $\operatorname{poly}$ many bits of randomness. Merlin then prepares a proof based on this shared randomness and sends it Arthur, who verifies the proof in combination with the random bits. In YES cases Arthur should accept with probability $2/3$ over the shared randomness, whereas he may accepts NO cases only with probability at most $1/3$.
The class $\mathsf{QAM}$ generalizes $\mathsf{AM}$ to the quantum setting by making Arthur a $\mathsf{BQP}$ verifier and the proof a quantum state. The shared randomness, however, is still classical. Here $\mathsf{QAM}$ differs from $\mathsf{qq\mhyphen QAM}$. In $\mathsf{qq\mhyphen QAM}$, the classical correlations between Arthur and Merlin (the shared randomness) are replaced by quantum correlations. That is, Arthur and Merlin share $\operatorname{poly}$ many EPR pairs instead of random bits. They could measure these in the computational basis to obtain shared randomness, showing that $\mathsf{QAM}\subseteq\mathsf{qq\mhyphen QAM}$, but are not required to do so. Formally, we define $\mathsf{qq\mhyphen QAM}$ as follows.
Definition 2.1 ($\mathsf{qq\mhyphen QAM}$-protocol).
A $\mathsf{qq\mhyphen QAM}$-protocol consists of a verifier and an all-powerful prover interacting. On an input $x$ of length $n$ they take the following actions.
  1. 1.
    The verifier prepares $\operatorname{poly}(n)$ EPR pairs across registers $A,B_{1}$. They send the $B_{1}$ register to the prover and keep the $A$ register themselves.
  2. 2.
    The prover performs any quantum operation on $B=B_{1}B_{2}$ and potentially on a private register $P$. They then send the $B$ register to the verifier who now has access to the proof state $\rho_{AB}$.
  3. 3.
    The verifier applies a $\operatorname{poly}(n)$-size quantum circuit $U_{x}$ to $\rho_{AB}$ and potentially to a $\operatorname{poly}(n)$ size workspace register $W$. They then measure the first qubit of $A$ in the computational basis, accept if the result is 1 and reject otherwise.
Definition 2.2 (quantum-quantum Quantum Arthur Merlin ($\mathsf{qq\mhyphen QAM}$) see also [KGN19]).
A promise problem $(L_{yes},L_{no})$ is in $\mathsf{qq\mhyphen QAM}(c,s)$ if there is some $\mathsf{qq\mhyphen QAM}$-protocol such that
  • $\forall x\in L_{yes}$, the verifier accepts with probability at least $c$.
  • $\forall x\in L_{no}$, the verifier accepts with probability at most $s$.
We define $\mathsf{qq\mhyphen QAM}=\mathsf{qq\mhyphen QAM}(\frac{2}{3},\frac{1}{3})$.
In [KGN19][Lemma 20] it is proven that $\mathsf{qq\mhyphen QAM}$ completeness and soundness parameters can be boosted using parallel repetition. That is, for $c,s$ satisfying $c-s=\frac{1}{\operatorname{poly}}$ and for any polynomial $p$, $\mathsf{qq\mhyphen QAM}(c,s)=\mathsf{qq\mhyphen QAM}=\mathsf{qq\mhyphen QAM}(1-2^{-p},2^{-p})$.
Note that by Lemma 2.14, the prover can make $\rho_{AB}$ be any state that satisfies $\rho_{A}=\widetilde{I}_{A}$. We can thus equivalently define $\mathsf{qq\mhyphen QAM}$ as $\mathsf{QMA}$ where the proof $\rho_{AB}$ is potentially mixed but promised to satisfy $\rho_{A}=\widetilde{I}_{A}$.
In [KGN19], it was shown that $\mathsf{qq\mhyphen QAM}$ has two natural complete problems.
Definition 2.3 (Close Image to Maximally Mixed problem (CIMM)).
The input to a $\mathsf{CIMM}$ instance consists of a quantum channel $\Phi$ and constants $a,b\in(0,1)$ such that $(1-a)^{2}>1-b^{2}$. The task is to distinguish the following two cases:
  • YES:
    $\exists\rho$ such that $\|\Phi(\rho)-\widetilde{I}\|_{1}\leq a$.
  • NO:
    $\forall\rho$, $\|\Phi(\rho)-\widetilde{I}\|_{1}\geq\beta$.
Theorem 2.4 ([KGN19]).
$\mathsf{CIMM}$ is complete for $\mathsf{qq\mhyphen QAM}$.
Definition 2.5 (Maximum Output Quantum Entropy Approximation ($\mathsf{MaxOutQEA}$)).
The input to a $\mathsf{MaxOutQEA}$ instance consists of a quantum channel $\Phi$ and a parameter $t\in\mathbb{R}$. The task is to distinguish the following two cases:
  • YES:
    $\exists\rho$ such that $S(\Phi(\rho))\geq t+1$.
  • NO:
    $\forall\rho$, $S(\Phi(\rho))\leq t-1$.
Theorem 2.6 ([KGN19, Theorem 3]).
$\mathsf{MaxOutQEA}$ is $\mathsf{qq\mhyphen QAM}$ complete.
Pure Product Isometry Output.
The following $\mathsf{QMA}(2)$-complete problem will be very useful for our purposes.
Definition 2.7 (Pure Product Isometry Output ($\mathsf{PPIO}$)).
The input to a $\mathsf{PPIO}$ instance is a quantum circuit implementing an isometry $U$ with a 2-partite [^9] The general definition allows $\ell$-partite output systems, but for our purposes the 2-partite case suffices. output system $AB$ and two parameters $a,b$. The task is to decide between:
  • YES:
    There is some isometry output that is close to product:
    $\exists\left|\psi\right\rangle,\left|\phi_{A}\right\rangle,\left|\phi_{B}\right\rangle\text{ such that }\|U\left(\lvert\psi\rangle\langle\psi\rvert_{A}\otimes\lvert 0^{n_{B}}\rangle\langle 0^{n_{B}}\rvert_{B}\right)U^{\dagger}-\lvert\phi_{A}\rangle\langle\phi_{A}\rvert\otimes\lvert\phi_{B}\rangle\langle\phi_{B}\rvert\|\leq a.$
  • NO:
    All isometry outputs are far from product:
    $\forall\left|\psi\right\rangle,\left|\phi_{A}\right\rangle,\left|\phi_{B}\right\rangle\text{, }\|U\left(\lvert\psi\rangle\langle\psi\rvert_{A}\otimes\lvert 0^{n_{B}}\rangle\langle 0^{n_{B}}\rvert_{B}\right)U^{\dagger}-\lvert\phi_{A}\rangle\langle\phi_{A}\rvert\otimes\lvert\phi_{B}\rangle\langle\phi_{B}\rvert\|\geq b.$
Theorem 2.8 ([GHMW13, Theorem 6.2]).
$\mathsf{PPIO}$ is $\mathsf{QMA}(2)$-complete, also for $a=2^{-\Omega(n)}$ and $b=2-2^{-\Omega(n)}$.
2.2 Useful definitions and lemmas
Quantum entropies.
In this paper we make use of two different quantum entropies: the Von Neumann entropy and the quantum min-entropy. The Von Neumann entropy is the quantum equivalent of the classical Shannon entropy and is defined as:
$S(\rho)=-\operatorname{Tr}\left(\rho\log\rho\right).$ (4)
Indeed, if $\rho$ is a classical probability distribution, the Shannon entropy and the Von Neumann entropy coincide.
The quantum min-entropy is defined as:
$H_{\infty}(\rho)=-\log\|\rho\|_{\infty}.$ (5)
Von Neumann entropy and min-entropy are related as follow:
$\forall\rho,\quad S(\rho)\geq H_{\infty}(\rho)$ (6)
Stinespring representation of channels.
By the Stinespring Dilation Theorem (see e.g. [Wil13][Section 5.2.2]), any quantum channel $\Phi\colon L(\mathcal{H}_{A})\to L(\mathcal{H}_{B})$ can be written as
$\Phi(\rho)=\operatorname{Tr}_{E}(V\rho V),$ (7)
where $V\colon L(\mathcal{H}_{A})\to L(\mathcal{H}_{B}\otimes\mathcal{H}_{E})$ is an isometry and $E$ is an auxiliary register with $\dim(\mathcal{H}_{E})\leq\dim(\mathcal{H}_{A})\dim(\mathcal{H}_{B})$. This way of writing $\Phi$ is often called the Stinespring representation of $\Phi$. On finite dimensional Hilbert spaces (the setting we consider), isometries can be extended to unitaries. It follows that, for every channel $\Phi$, there exists a unitary $U_{\Phi}$ acting on the $A,B$ and $E$ registers such that
$\Phi(\rho)=\operatorname{Tr}_{AE}\left(U_{\Phi}(\rho\otimes\lvert 0^{n_{B}}\rangle\langle 0^{n_{B}}\rvert_{B}\otimes\lvert 0^{n_{A}+n_{B}}\rangle\langle 0^{n_{A}+n_{B}}\rvert_{E})U_{\Phi}^{\dagger}\right).$ (8)
In this work, whenever we refer to the Stinespring representation of a channel, we refer to the unitary $U_{\Phi}$.
Quantum extractors.
One tool we will use are quantum extractors. Informally, these are quantum channels that “extract” the entropy from their input state, outputting a maximally mixed state if the input has sufficient entropy and something far from maximally mixed if it does not. More formally:
Definition 2.9 (Quantum extractors (see e.g. [BST08][section 5.1).
)] A quantum channel $T\colon L(\mathcal{H})\to L(\mathcal{H})$ is a $(k,d,\varepsilon)$-quantum extractor if it satisfies the following properties:
  1. 1.
    $T$ is $2^{d}$-regular. That is, there are unitary operators $\{U_{i}\}_{1\leq i\leq 2^{d}}$ such that $T$ can be written as
    $T(\rho)=\frac{1}{2^{d}}\sum_{i=1}^{2^{d}}U_{i}\rho U_{i}^{\dagger}$ (9)
  2. 2.
    For every $\rho$ with $H_{\infty}(\rho)\geq k$, one has $\|T(\rho)-\widetilde{I}\|_{1}\leq\varepsilon$.
A quantum extractor is called explicit if it can be implemented by a quantum circuit of size $\operatorname{poly}(\dim\mathcal{H})$.
Lemma 2.10 ([BST08][Corollary 5.3]).
For all $n,k,\varepsilon$, there exist explicit $(k,d,\varepsilon)$-quantum extractor $T\colon L((\mathbb{C}^{2})^{\otimes n})\to L((\mathbb{C}^{2})^{\otimes n})$ with $d=2(n-k+2\log\frac{1}{\varepsilon})+O(1)$, or $d=n-k+2\log\frac{1}{\varepsilon}+2\log n+O(1)$, depending on the construction used.
Swap test.
The SWAP test (see e.g. [BCWW01]) is a quantum subroutine that, when given as input a quantum state $\left|\phi\right\rangle\otimes\left|\psi\right\rangle$ accepts with probability $p_{\textup{acc}}=\frac{1}{2}+\frac{1}{2}|\left\langle\psi\middle|\phi\right\rangle|^{2}$. It can therefore be used to estimate the overlap of two pure states $\left|\phi\right\rangle$ and $\left|\psi\right\rangle$. Note that for pure states we have
$\|\phi-\psi\|_{1}$ $=\sqrt{1-|\left\langle\psi\middle|\phi\right\rangle|^{2}}$ (10)
$=\sqrt{2(1-p_{\textup{acc}})}.$ (11)
Therefore, the SWAP test can also be used to estimate the trace distance of two pure states.
Useful lemmas.
Lemma 2.11 (Gentle Measurement Lemma (see e.g. [Wil13][Lemma 9.4.1])).
Let $\rho$ be a quantum state and consider a measurement operator $0\preceq M\preceq I$ with $\operatorname{Tr}(M\rho)\geq 1-\varepsilon$. Then the post-measurement state $\rho_{post}=\frac{\sqrt{M}\rho\sqrt{M}}{\operatorname{Tr}(M\rho)}$ satisfies:
$\|\rho-\rho_{post}\|_{1}\leq 2\sqrt{\varepsilon}$ (12)
Theorem 2.12 (Uhlmann’s Theorem).
Let $\sigma$ and $\rho$ be two mixed states. Then
$F(\sigma,\rho)=\max_{\left|\phi_{\sigma}\right\rangle,\left|\phi_{\rho}\right\rangle}|\left\langle\phi_{\sigma}\middle|\phi_{\rho}\right\rangle|^{2}$ (13)
Lemma 2.13 (See [NC10][Exercise 2.81]).
Let $\left|\phi\right\rangle_{AB}$ and $\left|\psi\right\rangle_{AB}$ be two purifications of a mixed state $\rho_{A}$. Then there is a unitary $U_{B}$ acting only on the purifying register such that $I_{A}\otimes U_{B}\left|\phi\right\rangle=\left|\psi\right\rangle$.
Lemma 2.14.
Let $\rho_{A},\sigma_{A}$ be quantum states such that $\lVert\rho_{A}-\sigma_{A}\rVert_{\mathrm{tr}}\leq\varepsilon$ and let $\left|\phi_{\rho}\right\rangle_{AB},\left|\phi_{\sigma}\right\rangle_{AB}$ be purifications of $\rho,\sigma$. Then, there exists a unitary $U_{B}$ acting only on the purifying register $B$ such that $F\left(\lvert\phi_{\sigma}\rangle\langle\phi_{\sigma}\rvert,(I_{A}\otimes U_{B})\lvert\phi_{\rho}\rangle\langle\phi_{\rho}\rvert(I_{A}\otimes U_{B})\right)\geq(1-2\varepsilon)^{2}$.
Proof.
By the Fuchs-van de Graaf Inequality, $F(\rho,\sigma)\geq(1-2\varepsilon)^{2}$. By Uhlmann’s Theorem there are purifications $\left|\psi_{\rho}\right\rangle$ and $\left|\psi_{\sigma}\right\rangle$ with $|\left\langle\psi_{\sigma}\middle|\psi_{\rho}\right\rangle|^{2}\geq(1-2\varepsilon)^{2}$. By Lemma 2.13, there are unitaries $U_{\rho}$ and $U_{\phi}$ such that $(I\otimes U_{\rho})\left|\phi_{\rho}\right\rangle=\left|\psi_{\rho}\right\rangle$ and $(I\otimes U_{\sigma})\left|\phi_{\sigma}\right\rangle=\left|\psi_{\sigma}\right\rangle$. It is now clear that $U_{B}\coloneq U_{\sigma}^{\dagger}U_{\rho}$ has the claimed property. ∎
Lemma 2.15 (Fannes inequality (see e.g. [NC10][Theorem 11.6])).
Let $\rho$ and $\sigma$ be two quantum states on a $d$-dimensional Hilbert space and let $T=\|\rho-\sigma\|_{1}$. Then,
$|S(\rho)-S(\sigma)|\leq T\log d-T\log T,$ (14)
and
$|S(\rho)-S(\sigma)|\leq T\log d-\frac{1}{e\ln 2}$ (15)
Lemma 2.16 ([NC10, Theorem 11.10]).
Let $\rho=\sum_{x}p_{x}\rho_{x}$ be a quantum state, where the $p_{x}$ form a probability distribution. Then,
$S(\rho)\leq\sum_{x}p_{x}S(\rho_{x})+H(p),$ (16)
where $H(p)$ is the Shannon entropy of $p$.
Lemma 2.17 (Flattening Lemma (see e.g. [BST08][Lemma 5.6])).
Let $\rho$ be an $n$-qubit quantum state. Then, for all $\varepsilon>0$, there is a state $\sigma$ on $qn$ qubits such that $\|\rho^{\otimes q}-\sigma\|_{1}\leq 2\varepsilon$ and
$H_{\infty}(\sigma)\geq qS(\rho)-O\left(n+\log\frac{q}{\varepsilon}\right)\sqrt{q\log\frac{1}{\varepsilon}}.$ (17)
3 Main technical lemmas
3.1 Entropy verification in $\mathsf{qq\mhyphen QAM}$
Here we show how to verify entropy in $\mathsf{qq\mhyphen QAM}$. This approach has previously appeared in [KGN19][proof of Lemma 33] and is based on ideas from [BST08][Section 5]. We explicitly state the procedure as a lemma for completeness and to hopefully ease further use.
Let $A,B$ be registers of $n_{A},n_{B}$ qubits respectively, where $n_{A}+n_{B}=n$, and let $\tau=\tau(n)$ be an arbitrary function of $n$. Consider the registers $\mathcal{A}=A_{1}\dots A_{q}$ and $\mathcal{B}=B_{1}\dots B_{q}$ consisting of $q$ copies of the $A$ and $B$ registers. Fix an explicit $(k,d,\varepsilon)$-extractor $T\colon L(\mathcal{H}_{\mathcal{A}})\to L(\mathcal{H}_{\mathcal{A}})$ with $k=q\tau-O\left(n_{A}+\log\frac{q}{\varepsilon}\right)\sqrt{q\log\frac{1}{\varepsilon}}$, $d=qn_{A}-k+2\log\frac{qn_{A}}{\varepsilon}+O(1)$ and $q\leq\operatorname{poly}(n)$ and $\varepsilon\geq\frac{1}{\operatorname{poly}(n)}$ to be determined later. Such an extractor exists by Lemma 2.10. Let $U_{T}$ be an efficient unitary implementing $T$ acting on $\mathcal{A}$ and a $2qn_{A}$ qubit register $E$. That is, let $U_{T}$ be such that
$T(\rho)=\operatorname{Tr}_{E}(U_{T}(\rho\otimes\lvert 0^{2qn_{A}}\rangle\langle 0^{2qn_{A}}\rvert)U_{T}^{\dagger}).$ (18)
Now, consider the protocol, depending on $\tau$, $q$ and $\varepsilon$, that on input $\rho_{\mathcal{A}\mathcal{B}E}$ does the following:
  1. 1.
    Apply $U_{T}^{\dagger}\otimes I_{\mathcal{B}}$ to to the input state.
  2. 2.
    Perform a projective measurement $\{\Pi,I-\Pi\}$ where $\Pi=I_{\mathcal{A}}\otimes I_{\mathcal{B}}\otimes\lvert 0^{2qn_{A}}\rangle\langle 0^{2qn_{A}}\rvert_{E}$. Reject if the outcome is not $\Pi$.
  3. 3.
    Output the resulting state $\sigma_{\mathcal{A}\mathcal{B}}$.
Note that the $\mathcal{B}$ register is not acted on at all. Its sole purpose is to allow for correlations with a system whose entropy is not checked.
Lemma 3.1 (Entropy verification in $\mathsf{qq\mhyphen QAM}$).
For all $\tau$ and all $\delta,\delta^{\prime}\geq\frac{1}{\operatorname{poly}(n)}$ there are $c,s,\varepsilon,q$ with $c-s\geq\frac{1}{\operatorname{poly}(n)}$, $\varepsilon\geq\frac{1}{\operatorname{poly}(n)}$ and $q\leq\operatorname{poly}(n)$ such that the entropy verification protocol above satisfies the following properties:
  1. 1.
    For any state $\rho_{AB}$ with $S(\rho_{A})\geq\tau$ there exists some state $\chi_{\mathcal{A}\mathcal{B}E}$ with $\chi_{\mathcal{A}}=\widetilde{I}_{\mathcal{A}}$ that is accepted by the protocol with probability at least $c$. The output state $\sigma_{\mathcal{A}\mathcal{B}}$ in case of acceptance satisfies
    $\left\|\rho_{AB}^{\otimes q}-\sigma_{\mathcal{A}\mathcal{B}}\right\|_{1}\leq\delta.$ (19)
    Furthermore, the average output state $\tilde{\sigma}_{AB}\coloneq\frac{1}{q}\sum\sigma_{A_{i}B_{i}}$ has $\|\rho_{AB}-\tilde{\sigma}_{AB}\|_{1}\leq\delta$.
  2. 2.
    Let $\chi_{\mathcal{A}\mathcal{B}E}$ be an arbitrary state satisfying $\chi_{\mathcal{A}}=\widetilde{I}_{\mathcal{A}}$ and suppose that on input $\chi$, the measurement accepts with probability at least $s$. Then, the output state in case of acceptance $\sigma_{\mathcal{A},\mathcal{B}}$ satisfies $S(\sigma_{\mathcal{A}})\geq q(\tau-\delta^{\prime})$ and the average output state $\tilde{\sigma}_{AB}$ has $S(\tilde{\sigma}_{A})\geq\tau-\delta^{\prime}$.
Proof.
For the first property, let $\rho_{AB}$ satisfy $S(\rho_{A})\geq\tau$ and consider
$\chi_{\mathcal{A}\mathcal{B}E}=(U_{T}\otimes I_{\mathcal{B}})\left(\rho_{A_{1}B_{1}}\otimes\dots\otimes\rho_{A_{q}B_{q}}\otimes\lvert 0^{2qn_{A}}\rangle\langle 0^{2qn_{A}}\rvert_{E}\right)(U_{T}^{\dagger}\otimes I_{\mathcal{B}}).$ (20)
Note that by definition of $U_{T}$, $\operatorname{Tr}_{\mathcal{B}E}(\chi)=T(\rho_{A}^{\otimes q})$. As the entropy of $\rho_{A}$ is high by assumption, the Flattening Lemma guarantees the existence of a state $\eta_{\mathcal{A}}$ with $\|\rho_{A}^{\otimes q}-\eta_{\mathcal{A}}\|_{1}\leq 2\varepsilon$ and
$H_{\infty}(\eta_{\mathcal{A}})\geq q\tau-O\left(n_{A}+\log\frac{q}{\varepsilon}\right)\sqrt{q\log\frac{1}{\varepsilon}}=k.$ (21)
By the second property of Definition 2.9 we have:
$\|\operatorname{Tr}_{\mathcal{B}E}(\chi)-\widetilde{I}_{\mathcal{A}}\|_{1}$ $=\|T(\rho_{A}^{\otimes q})-\widetilde{I}_{\mathcal{A}}\|_{1}$ (22)
$\leq\|T(\rho^{\otimes q})-T(\eta_{\mathcal{A}})\|_{1}+\|T(\sigma_{\mathcal{A}})-\widetilde{I}_{\mathcal{A}}\|_{1}$ (23)
$\leq 2\varepsilon+\varepsilon,$ (24)
where we used that trace distance does not increase under quantum channels. By Lemma 2.14 there exists a state $\tilde{\chi}_{\mathcal{A}\mathcal{B}E}$ with $\operatorname{Tr}_{\mathcal{B}E}(\tilde{\chi}_{\mathcal{A}\mathcal{B}E})=\widetilde{I}_{\mathcal{A}}$ and $F(\chi,\tilde{\chi})\geq(1-6\varepsilon)^{2}\geq 1-12\varepsilon^{2}$.
Using that fidelity is invariant under unitaries and the Fuchs-van de Graaf Inequality, we see that (dropping the identity applied to the $\mathcal{B}$ register for legibility)
$\begin{split}\|U_{T}^{\dagger}\tilde{\chi}_{\mathcal{A}\mathcal{B}E}U_{T}-\rho_{AB}^{\otimes q}\otimes\lvert 0^{2qn_{A}}\rangle\langle 0^{2qn_{A}}\rvert_{E}\|_{1}&\leq\|U_{T}^{\dagger}\tilde{\chi}U_{T}-U_{T}^{\dagger}\chi U_{T}\|_{1}\\ &\qquad+\|U_{T}^{\dagger}\chi U_{T}-\rho^{\otimes q}\otimes\lvert 0^{2qn_{A}}\rangle\langle 0^{2qn_{A}}\rvert\|_{1}\end{split}$ (25)
$=\|\tilde{\chi}-\chi\|_{1}+0$ (26)
$\leq 4\sqrt{3\varepsilon}.$ (27)
It follows that the measurement accepts with probability at least $c\coloneq 1-4\sqrt{3\varepsilon}$. By the Gentle Measurement Lemma, the post-measurement state
$\sigma_{\mathcal{A}\mathcal{B}E}=\frac{\Pi U_{T}^{\dagger}\tilde{\chi}_{\mathcal{A}\mathcal{B}E}U_{T}\Pi}{\operatorname{Tr}\left(\Pi U_{T}^{\dagger}\tilde{\chi}_{\mathcal{A}\mathcal{B}E}U_{T}\right)}$ (28)
satisfies
$\|\sigma_{\mathcal{A}\mathcal{B}}-\rho_{AB}^{\otimes q}\|_{1}\leq 4\sqrt[4]{3\varepsilon},$ (29)
which immediately implies $\|\tilde{\sigma}_{AB}-\rho_{AB}\|_{1}\leq 4\sqrt[4]{3\varepsilon}$. Choosing $\varepsilon$ such that $4\sqrt[4]{3\varepsilon}\leq\delta$ we see that $\tilde{\chi}$ is as claimed.
We now turn to the second property. Let $\chi_{\mathcal{A}\mathcal{B}E}$ be a state with $\chi_{\mathcal{A}}=\widetilde{I}_{\mathcal{A}}$ and assume that
$\operatorname{Tr}\left(\Pi\left(U_{T}^{\dagger}\otimes I_{\mathcal{B}}\right)\chi_{\mathcal{A}\mathcal{B}E}\left(U_{T}\otimes I_{\mathcal{B}}\right)\right)\geq s$ (30)
By the Gentle Measurement Lemma, $U_{T}^{\dagger}\chi U_{T}$ is close to the post-measurement state $\sigma_{\mathcal{A}\mathcal{B}}$:
$\left\|(U_{T}^{\dagger}\otimes I_{\mathcal{B}})\chi_{\mathcal{A}\mathcal{B}E}(U_{T}\otimes I_{B})-\sigma_{\mathcal{A}\mathcal{B}}\otimes\lvert 0^{2qn_{A}}\rangle\langle 0^{2qn_{A}}\rvert_{E}\right\|_{1}\leq 2\sqrt{1-s}.$ (31)
Now, consider $T(\sigma)=\operatorname{Tr}_{\mathcal{B}E}(U_{T}(\sigma\otimes\lvert 0^{2qn_{A}}\rangle\langle 0^{2qn_{A}}\rvert)U_{T}^{\dagger})$ (omitting the $I_{\mathcal{B}}$). Using that trace distance is non-increasing under partial trace and invariant under unitaries, we have:
$\|T(\sigma)-\widetilde{I}_{\mathcal{A}}\|_{1}$ $\leq\|T(\sigma)-\chi_{\mathcal{A}}\|_{1}+\|\chi_{\mathcal{A}}-\widetilde{I}_{\mathcal{A}}\|_{1}$ (32)
$\leq\|U_{T}(\sigma_{\mathcal{A}}\otimes\lvert 0^{2qn_{A}}\rangle\langle 0^{2qn_{A}}\rvert_{E})U_{T}^{\dagger}-\chi_{\mathcal{A}E}\|_{1}+0$ (33)
$\leq\|\sigma_{\mathcal{A}}\otimes\lvert 0^{2qn_{A}}\rangle\langle 0^{2qn_{A}}\rvert_{E}-U_{T}^{\dagger}\chi_{\mathcal{A}E}U_{T}\|_{1}$ (34)
$\leq 2\sqrt{1-s}.$ (35)
By the Fannes Inequality:
$S(T(\sigma_{\mathcal{A}}))\geq qn_{A}-2\sqrt{1-s}\left(qn_{A}-\log(2\sqrt{1-s})\right).$ (36)
Because $T$ is $2^{d}$-regular, it can increase the entropy by at most $d$. It follows that:
$S(\sigma_{\mathcal{A}})\geq qn_{A}-2\sqrt{1-s}\left(qn_{A}-\log(2\sqrt{1-s})\right)-d.$ (37)
Recall that $k=q\tau-O\left(n_{A}+\log\frac{q}{\varepsilon}\right)\sqrt{q\log\frac{1}{\varepsilon}}$ and that $d=qn_{A}-k+2\log\frac{qn_{A}}{\varepsilon}+O(1)$. Therefore $d=q(n_{A}-\tau)+O\left(n_{A}+\log\frac{q}{\varepsilon}\right)\sqrt{q\log\frac{1}{\varepsilon}}$ and we have:
$S(\sigma)\geq q\left(\tau-2\sqrt{1-s}\left(n_{A}-\frac{1}{q}\log(2\sqrt{1-s})\right)-O(n_{A}+\log\frac{q}{\varepsilon})\sqrt{\frac{1}{q}\log\frac{1}{\varepsilon}}\right).$ (38)
We would like $S(\sigma)\geq q(\tau-\delta^{\prime})$, which requires
$2\sqrt{1-s}\left(n_{A}-\frac{1}{q}\log(2\sqrt{1-s})\right)+O(n_{A}+\log\frac{q}{\varepsilon})\sqrt{\frac{1}{q}\log\frac{1}{\varepsilon}}\leq\delta^{\prime}.$ (39)
We now need to set $\varepsilon,q$ and $s$ such that $c-s\geq\frac{1}{\operatorname{poly}(n)}$, $4\sqrt[4]{3\varepsilon}\leq\delta$ and Eq. 39 are satisfied (recall that we previously set $c=1-4\sqrt{3\varepsilon})$. This can indeed be achieved by first setting $s$ such that $2\sqrt{1-s}n\leq\frac{\delta^{\prime}}{4}$. Note that an $s$ with $1-s\geq\frac{1}{\operatorname{poly}(n)}$ suffices. We then set $\varepsilon$ such that $c-s=1-s-4\sqrt{3\varepsilon}\geq\frac{1}{\operatorname{poly}(n)}$ and $4\sqrt[4]{3\varepsilon}\leq\delta$, for which $\varepsilon\geq\frac{1}{\operatorname{poly}(n)}$ works. Finally, we set $q$ to a large enough polynomial in $n$ such that Eq. 39 holds.
Now, consider the average output state $\tilde{\sigma}_{AB}=\frac{1}{q}\sum_{i=1}^{q}\sigma_{A_{i}B_{i}}$. By concavity and subadditivity of Von Neumann entropy we have
$S(\tilde{\sigma}_{A})$ $=S\left(\frac{1}{q}\sum_{i=1}^{q}\sigma_{A_{i}}\right)$ (40)
$\geq\frac{1}{q}\sum_{i=1}^{q}S(\sigma_{A_{i}})$ (41)
$\geq\frac{1}{q}S(\sigma_{\mathcal{A}})$ (42)
$\geq\tau-\delta^{\prime},$ (43)
completing the proof. ∎
3.2 Channel-to-Hamiltonian construction
The main tool for our hardness proofs is a construction that allows embedding the range of a quantum channel in the ground space of local Hamiltonian.
Lemma 3.2 (Channel-to-Hamiltonian construction).
Let $\Phi\colon L(\mathcal{H}_{A})\to L(\mathcal{H}_{B})$ be a quantum channel and let $n_{A},n_{B}$ be the number of qubits of $\mathcal{H}_{A}$ and $\mathcal{H}_{B}$ respectively. Let $E$ be a register of $n_{A}+n_{B}$ qubits and consider a unitary $U_{\Phi}=V_{T}\dots V_{1}$ acting on $ABE$ implementing the Stinespring representation of $\Phi$:
$\Phi(\rho)=\operatorname{Tr}_{AE}\left(U_{\Phi}(\rho_{A}\otimes\lvert 0^{n_{B}}\rangle\langle 0^{n_{B}}\rvert_{B}\otimes\lvert 0^{n_{A}+n_{B}}\rangle\langle 0^{n_{A}+n_{B}}\rvert_{E})U_{\Phi}^{\dagger}\right).$ (44)
Then, there exists some $5$-local Hamiltonian $H_{\Phi}$, depending on a parameter $L$ and acting on the registers $A,B,E$ and a $T+L$-qubit clock register $C$ such that
  1. 1.
    For any $\sigma_{A}$, there exists some $\rho_{ABCE}$ (pure if $\sigma$ is pure) such that $\operatorname{Tr}(H\rho)=0$ and $\|\rho_{B}-\Phi(\sigma)\|_{1}\leq\frac{2T}{T+L+1}$.
  2. 2.
    If $\rho_{ABCE}$ is such that $\operatorname{Tr}(H\rho)\leq\beta$, then there exists some state $\sigma_{A}$ (pure if $\rho$ is pure) such that $\|\rho_{B}-\Phi(\sigma)\|_{1}\leq 2\sqrt{\frac{\beta}{\Delta}}+\frac{2T}{T+L+1}$. Here $\Delta$ is the spectral gap of $H_{\Phi}$, which satisfies $\Delta=\Omega\left((T+L+1)^{-3}\right)$.
Proof.
We will use Kitaev’s circuit-to-Hamiltonian construction applied to the unitary $U_{\Phi}$ with two modification. We use no output penalty and post idle for $L$ time steps. The Hamiltonian $H_{\Phi}$ will be given by
$H_{\Phi}$ $=H_{\textup{in}}+H_{\textup{prop}}+H_{\textup{clock}}$ (45)
$H_{\textup{in}}$ $=\left(\sum_{i=1}^{n_{B}}(I_{B_{i}}-\lvert 1\rangle\langle 1\rvert_{B_{i}})+\sum_{j=1}^{n_{A}+n_{B}}(I_{E_{i}}-\lvert 1\rangle\langle 1\rvert_{E_{i}})\right)\otimes\lvert 0\rangle\langle 0\rvert_{C}$ (46)
$H_{\textup{prop}}$ $=\sum_{t=0}^{T+L-1}-V_{t+1}\otimes\lvert t+1\rangle\langle t\rvert_{C}-V_{t+1}^{\dagger}\otimes\lvert t\rangle\langle t+1\rvert_{C}+I\otimes\lvert t\rangle\langle t\rvert_{C}+I\otimes\lvert t+1\rangle\langle t+1\rvert_{C}$ (47)
$H_{\textup{clock}}$ $=I_{ABE}\otimes\sum_{i=1}^{T+L-1}\lvert 01\rangle\langle 01\rvert_{C_{i}C_{i+1}},$ (48)
where $V_{t}=I$ for $t>T$.
The ground state of $H_{\Phi}$ will now consist of history states of the form:
$\left|\psi_{\textup{hist}}\right\rangle$ $=\frac{1}{\sqrt{T+L+1}}\sum_{t=0}^{T+L}V_{t}\dots V_{1}\left|\psi\right\rangle_{A}\left|0^{n_{B}}\right\rangle_{B}\left|0^{n_{A}+n_{B}}\right\rangle_{E}\left|t\right\rangle_{C}$ (49)
$\begin{split}&=\frac{1}{\sqrt{T+L+1}}\sum_{t=0}^{T-1}V_{t}\dots V_{1}\left|\psi\right\rangle_{A}\left|0^{n}\right\rangle_{B}\left|0^{n_{A}+n_{B}}\right\rangle_{E}\left|t\right\rangle_{C}\\ &\qquad+U_{\Phi}\left|\psi\right\rangle_{A}\left|0^{n}\right\rangle_{B}\left|0^{n_{A}+n_{B}}\right\rangle_{E}\left(\frac{1}{\sqrt{T+L+1}}\sum_{t=T}^{T+L}\left|t\right\rangle_{C}\right),\end{split}$ (50)
where $\left|t\right\rangle_{C}=\left|1^{t}0^{T+L-1-t}\right\rangle$ encodes the time step $t$ in unary. As we use no output penalty, any history state will have zero energy and all zero-energy states will be history states. It is shown in [GK12][Lemma 23] that Kitaev Hamiltonians without output penalty, and in particular $H_{\Phi}$, have spectral gap $\Delta\geq\Omega((T+L+1)^{-3})$. [^10] Note that our use of $\Delta$ differs from that of [GK12]. We use it for the spectral gap of $H_{\Phi}$ whereas [GK12] use it for a scaling constant.
To prove property 1, we start by considering pure states. Let $\left|\psi\right\rangle_{A}$ be an arbitrary pure state. The reduced state on the $B$ register of $\left|\psi_{hist}\right\rangle$ the history state on input $\left|\psi\right\rangle_{A}$ is given by [^11] Note that there are no cross-terms as all terms in the sum are orthogonal in the $C$ register.
$\eta_{\psi}\coloneq\operatorname{Tr}_{ACE}\left(\lvert\psi_{hist}\rangle\langle\psi_{hist}\rvert_{ABCE}\right)=\frac{L+1}{T+L+1}\Phi(\psi)+\frac{1}{T+L+1}\sum_{t=0}^{T-1}\rho^{\prime}_{t,\psi},$ (51)
where $\rho^{\prime}_{t,\psi}=\operatorname{Tr}_{ACE}\left(V_{t}\dots V_{1}\left|\psi\right\rangle_{A}\left|0^{n_{B}}\right\rangle_{B}\left|0^{n_{A}+n_{B}}\right\rangle_{E}\left|t\right\rangle_{C}\right)$ are error terms.
Now consider an arbitrary mixed state $\rho_{A}=\sum_{x}a_{x}\lvert\psi_{x}\rangle\langle\psi_{x}\rvert_{A}$. Write $\left|\psi_{hist,x}\right\rangle$ for the history state on input $\left|\psi_{x}\right\rangle$. By linearity of $\Phi$ and the partial trace we have:
[…]
$=\frac{L+1}{T+L+1}\sum_{x}a_{x}\Phi(\psi_{x})+\frac{1}{T+L+1}\sum_{x}a_{x}\sum_{t=0}^{T-1}\rho^{\prime}_{t,x}$ (53)
$=\frac{L+1}{T+L+1}\Phi(\rho)+\frac{1}{T+L+1}\sum_{t=0}^{T-1}\rho^{\prime\prime}_{t}.$ (54)
Here the $\rho^{\prime\prime}_{t}=\sum_{x}a_{x}\rho^{\prime}_{t,x}$ are again error terms. The trace distance between $\eta_{\rho}$ and $\Phi(\rho)$ can be upper bounded as
$\|\Phi(\rho)-\eta_{\rho}\|_{1}$ $=\left\|\frac{T}{T+L+1}\Phi(\rho)-\frac{1}{T+L+1}\sum_{t=0}^{T-1}\rho^{\prime\prime}_{t}\right\|_{1}$ (55)
$=\frac{T}{T+L+1}\left\|\Phi(\rho)-\frac{1}{T}\sum_{t=0}^{T-1}\rho^{\prime\prime}_{t}\right\|_{1}$ (56)
$\leq\frac{2T}{T+L+1}.$ (57)
To prove the second property, let $\rho_{ABCE}$ be some state with $\operatorname{Tr}(H\rho)\leq\beta$ and let $\Pi$ be the projector on the span of the history states, or equivalently, on the zero-energy space of $H_{\Phi}$. Note that $\beta\geq\operatorname{Tr}(H\rho)\geq\Delta\operatorname{Tr}((I-\Pi)\rho)=\Delta\left(1-\operatorname{Tr}(\Pi\rho)\right)$. Hence $\operatorname{Tr}(\Pi\rho)\geq 1-\frac{\beta}{\Delta}$. By the Gentle Measurement Lemma, the state $\eta_{ABCE}=\frac{\Pi\rho\Pi}{\operatorname{Tr}(\Pi\rho)}$ now satisfies $\|\rho-\eta\|_{1}\leq 2\sqrt{\frac{\beta}{\Delta}}$. Furthermore, $\eta$ is a mixture of history states. That is, there are states $\left|\psi_{x}\right\rangle_{A}$ such that $\eta=\sum_{x}a_{x}\lvert\psi_{hist,x}\rangle\langle\psi_{hist,x}\rvert$ where $\left|\psi_{hist,x}\right\rangle$ is the history state on input $\left|\psi_{x}\right\rangle$. Note that if $\rho$ is pure, then $\eta$ is pure and a sum over a single item suffices. As in Eq. 54 we have
$\eta_{B}=\frac{L+1}{T+L+1}\Phi(\sigma)+\frac{1}{T+L+1}\sum_{t=0}^{T-1}\sigma^{\prime}_{t},$ (58)
where $\sigma=\sum_{x}a_{x}\lvert\psi_{x}\rangle\langle\psi_{x}\rvert$ is in particular pure if $\rho$ is pure, and the $\sigma^{\prime}_{t}$ are error terms. The distance between $\rho_{B}$ and $\Phi(\sigma)$ is given by:
$\|\rho_{B}-\Phi(\sigma)\|_{1}$ $\leq\|\rho_{B}-\eta_{B}\|_{1}+\|\eta_{B}-\Phi(\sigma)\|_{1}$ (59)
$\leq 2\sqrt{\frac{\beta}{\Delta}}+\frac{2T}{T+L+1},$ (60)
completing the proof. ∎
4 The High Entropy Ground State problem
See 1.2 Here we show completeness for general Hamiltonians. Hardness for physical Hamiltonians is postponed to Section 8. We begin by showing containment.
Lemma 4.1.
For any constant $k$, $\mathsf{HELES}\in\mathsf{qq\mhyphen QAM}$.
Proof.
We will show the slightly stronger statement that $\mathsf{HELES}\in\mathsf{qq\mhyphen QAM}$, even if we replace the pure states in the definition of $\mathsf{HELES}$ with mixed states. [^12] Note that the YES and NO cases of the pure state variant are subsets of the YES and NO cases of the mixed state variant. Hence the mixed state variant is at least as hard as the pure state variant. In fact, our results show that they are equally hard.
Let $(H_{AB},\alpha,\beta,\mathfrak{s},\mathfrak{t})$ be a $\mathsf{HELES}$ instance. We use the entropy verification protocol from Lemma 3.1 to ensure that the prover sends a state with high entropy on the $A$ register. Let $q,\mathcal{A},\mathcal{B}$ and $E$ be as in Lemma 3.1. We split the $E=E_{1}E_{2}$ register in 2 halves of $qn_{A}$ qubits each. The $\mathcal{A}$ register will contain the verifiers halves of the EPR states, and the provers halves will be in the $E_{1}$ register. The prover will act on the $\mathcal{B}$ and $E$ registers and send them to the verifier who then has access to the entire proof state $\chi_{\mathcal{A}\mathcal{B}E}$.
The protocol for verifying $\mathsf{HELES}$ in $\mathsf{qq\mhyphen QAM}$ looks like this:
  1. 1.
    The verifier prepares EPR pairs across the $\mathcal{A}$ and $E_{1}$ registers. They then send the $E_{1}$ register to the prover.
  2. 2.
    The prover now acts on the $\mathcal{B}$ and $E$ registers, and potentially their own private register $\mathcal{P}$. They send the $\mathcal{B}$ and $E$ registers to the verifier, who then has access to the entire proof state $\chi_{\mathcal{A}\mathcal{B}E}$, which is potentially entangled with the $\mathcal{P}$ register.
  3. 3.
    The verifier runs the entropy verification protocol with $\tau=\mathfrak{s},\delta=\frac{\beta-\alpha}{4\|H\|_{\infty}}$ and $\delta^{\prime}<\mathfrak{s}-\mathfrak{t}$. They reject if the protocol rejects.
  4. 4.
    If they have not rejected yet in the previous step, they proceed by measuring the energy of the post-measurement state $\sigma_{\mathcal{A}\mathcal{B}}$. For each $A_{i}B_{i}$ pair, they pick a local term $H_{j}$ from the Hamiltonian $H$ uniformly at random and measure $\operatorname{Tr}(H_{j}\sigma_{A_{i}B_{i}})$. Taking the average over all $q$ pairs $A_{i}B_{i}$, they obtain an estimate $\tilde{E}$ of $\operatorname{Tr}(H\tilde{\sigma}_{AB})$. The verifier now accepts if $\tilde{E}<\frac{\alpha+\beta}{2}$ and rejects if it is bigger or equal to this value.
Note that because the prover could not act on the $\mathcal{A}$ register, the verifier is sure that $\chi_{\mathcal{A}}=\widetilde{I}_{\mathcal{A}}$. Furthermore, by Lemma 2.13, the prover can turn their halves of the EPR pairs into any state $\chi_{\mathcal{A}\mathcal{B}E}$ satisfying $\chi_{\mathcal{A}}=\widetilde{I}_{\mathcal{A}}$, without acting on the $\mathcal{A}$ register.
For completeness, let $\rho_{AB}$ be such that $\operatorname{Tr}(H\rho)\leq\alpha$ and $S(\rho_{A})\geq\mathfrak{s}$. The prover prepares the proof state $\chi_{\mathcal{A}\mathcal{B}E}$ such that the entropy verification protocol accepts with probability at least $c$, in which case the output state $\sigma_{\mathcal{A}\mathcal{B}}$ satisfies $\|\sigma_{\mathcal{A}\mathcal{B}}-\rho_{AB}^{\otimes q}\|_{1}\leq\delta$. This is possible by property 1 of Lemma 3.1.
Now consider the energy verification procedure. We have
$\frac{1}{q}\sum_{i=1}^{q}\mathbb{E}_{j}\left(\operatorname{Tr}(H_{j}\sigma_{A_{i}B_{i}})\right)$ $=\frac{1}{qm}\sum_{i=q}^{q}\sum_{j=1}^{m}\operatorname{Tr}(H_{j}\sigma_{A_{i}B_{j}})$ (61)
[…]
$\operatorname{Tr}(H\sigma)$ $=\operatorname{Tr}(H\rho)+\operatorname{Tr}(H(\sigma-\rho))$ (64)
$\leq\alpha+\|H(\sigma-\rho)\|_{1}$ (65)
$\leq\alpha+\|H\|_{\infty}\|\sigma-\rho\|_{1}$ (66)
$\leq\alpha+\frac{\beta-\alpha}{4}.$ (67)
By a Hoeffding bound, the probability that $\tilde{E}\geq\frac{\beta-\alpha}{2}$ is exponentially small in $q=\Omega(n)$. If desired, this can be decreased ever further with parallel repetition. In total, the acceptance probability is at least $c-2^{-\Omega(n)}$ in the YES case.
For soundness, suppose that any $\rho_{AB}$ with $S(\rho_{A})>\mathfrak{t}$ has $\operatorname{Tr}(H\rho)\geq\beta$. Let $\chi_{\mathcal{A}\mathcal{B}E}$ be the full proof state after step 2. By the second property of Lemma 3.1 we can distinguish two cases: either the entropy verification protocol accepts $\chi$ with probability less than $s$, or the average output state $\tilde{\sigma}$ has $S(\tilde{\sigma}_{A})\geq\tau-\delta^{\prime}>\mathfrak{t}$, and hence $\operatorname{Tr}(H\tilde{\sigma})\geq\beta$. In the first case, the total acceptance probability is at most $s$ which has an inverse polynomial gap with $c-2^{-\operatorname{poly}(n)}$ already. In the second case, the probability that $\tilde{E}<\frac{\alpha+\beta}{2}$ is exponentially small in $q$ by a Hoeffding bound. In either case, the acceptance probability in the NO case is hence at least inverse polynomially smaller than the acceptance probability in the YES case. Using parallel repetition this can be boosted to $\frac{1}{3},\frac{2}{3}$ respectively.
Lemma 4.2.
For $k\geq 5$, $\mathsf{HELES}$ is $\mathsf{qq\mhyphen QAM}$-hard.
Proof.
We show a poly-time many-one reduction from $\mathsf{MaxOutQEA}$. Let $(\Phi,\tau)$ be a $\mathsf{MaxOutQEA}$ instance. Without loss of generality we can assume that for all $\sigma$, there is a pure state $\left|\psi\right\rangle$ with $\Phi(\sigma)=\Phi(\psi)$. [^13] To see this, modify $\Phi\colon L(\mathcal{H}_{A})\to L(\mathcal{H}_{B})$ by adding another copy of the input register that is immediately traced out. Taking $\left|\psi\right\rangle$ to be some purification of $\sigma$ then suffices. Let $H_{\Phi}$ be the Hamiltonian acting on $A,B,C,E$ resulting from Lemma 3.2, with $L$ to be determined later. The output of the reduction will be $(H_{\Phi},\alpha,\beta,s,t)$, where $\alpha,\beta,s$ and $t$ will also be determined later.
For completeness, suppose $(\Phi,\tau)$ is a YES instance. That is, suppose there is some $\sigma_{A}$ with $S\left(\Phi(\sigma)\right)\geq\tau+1$. Let $\left|\psi\right\rangle_{A}$ be a pure state such that $\Phi(\psi)=\Phi(\sigma)$. By property 1 of Lemma 3.2 there exists a zero-energy pure state $\left|\phi\right\rangle_{ABCE}$ such that $\|\phi_{B}-\Phi(\psi)\|\leq\frac{2T}{T+L+1}$. The Fannes Inequality now gives
$S(\phi_{B})$ $\geq S\left(\Phi(\psi)\right)-\frac{2T}{T+L+1}\left(n_{B}-\log\frac{2T}{T+L+1}\right)$ (68)
Setting $\alpha=0$, $s=\tau+\frac{3}{4}$ and choosing $L$ such that
$\frac{2T}{T+L+1}\left(n_{B}-\log\frac{2T}{T+L+1}\right)\leq\frac{1}{4},$ (69)
we see that $\left|\phi\right\rangle_{ABCE}$ satisfies $S(\phi_{B})\geq s$ and $\left\langle\phi\right|H\left|\phi\right\rangle\leq\alpha$ as desired.
For soundness, suppose that for all $\sigma_{A}$, $S\left(\Phi(\sigma)\right)\leq\tau$ and let $\left|\psi\right\rangle_{ABCE}$ be such that $\left\langle\psi\right|H\left|\psi\right\rangle\leq\beta$ ($\beta$ will be determined later). By property 2 of Lemma 3.2, there is a pure state $\left|\phi\right\rangle_{A}$ such that
$T\coloneq\|\psi_{B}-\Phi(\phi_{A})\|_{1}\leq 2\sqrt{\frac{\beta}{\Delta}}+\frac{2T}{T+L+1}.$ (70)
From the Fannes Inequality it follows that
$S(\psi_{B})$ $\leq S(\Phi(\phi_{A}))+T(n_{B}-\log T)$ (71)
$\leq\tau+\left(\sqrt{\frac{\beta}{\Delta}}+\frac{2T}{T+L+1}\right)\left(n_{B}-\log\left(\sqrt{\frac{\beta}{\Delta}}+\frac{2T}{T+L+1}\right)\right).$ (72)
Choosing $t=\tau+\frac{1}{4}$, $\beta=\frac{\Delta}{L}=\Omega(L^{-4})$ and then setting $L$ large enough so that
$\left(\sqrt{\frac{\beta}{\Delta}}+\frac{2T}{T+L+1}\right)\left(n_{B}-\log\left(\sqrt{\frac{\beta}{\Delta}}+\frac{2T}{T+L+1}\right)\right)\leq\frac{1}{4}$ (73)
we see that $S(\psi_{B})\leq t$ as desired. Note that Eqs. 69 and 73 can be satisfied by an $L$ that is polynomially large in $n=n_{A}+n_{B}$. ∎
5 Free energy and Gibbs states
Definition 5.1 (Free energy).
The free energy of a Hamiltonian at inverse temperature $\beta$ is defined as:
$\mathcal{F}(H)=\min_{\rho}f(\rho)\quad\text{where}\quad f(\rho)=\operatorname{Tr}(H\rho)-\frac{S(\rho)}{\beta}.$ (74)
Equivalently, the free energy is also equal to
$\mathcal{F}(H)=\frac{\log\mathcal{Z}}{\beta}.$ (75)
Here $\mathcal{Z}=\operatorname{Tr}(e^{-\beta H})$ is also known as the partition function.
The state that minimizes the free energy is known as the Gibbs state and is given by $\rho_{\beta}=\frac{1}{\mathcal{Z}}e^{-\beta H}$.
See 1.3 See 1.4
Proof.
Let $(H,\beta,a,b)$ be an $\mathsf{FEA}$ instance, where $H=\sum_{j=1}^{m}H_{j}$ is a $k$-local Hamiltonian on an $n$-qubit register $A$. Let $\mathcal{A}=A_{1}\dots A_{q}$ and $E$ be as in Lemma 3.1. We do not use the $\mathcal{B}$ register and split $E=E_{1}E_{2}$ into two registers of $qn$ qubits each.
The verifiers actions will be similar as that in the proof of $\mathsf{HELES}\in\mathsf{qq\mhyphen QAM}$ (Lemma 4.1) and consists of the following steps.
  1. 1.
    Verifier prepares EPR pairs across the $\mathcal{A}$ and $E_{1}$ registers. They then send the $E_{1}$ register to the prover.
  2. 2.
    The prover now acts on the $E$ register, and potentially their own private register $\mathcal{P}$. They send the $E$ register to the verifier, who then has access to the entire proof state $\chi_{\mathcal{A}E}$. Furthermore, they send a classical value $\tilde{S}$ to the verifier.
  3. 3.
    The verifier runs the entropy verification protocol with $\delta=\frac{b-a}{4\|H\|_{\infty}},\delta^{\prime}<(1+\frac{1}{\beta})^{-1}\frac{b-a}{4}$ and $\tau=\tilde{S}$. They reject if the protocol rejects.
  4. 4.
[…]
For completeness, let $\rho_{\beta}=\frac{1}{\mathcal{Z}}e^{-\beta H}$ be the Gibbs state at inverse temperature $\beta$. In the YES case this satisfies $f(\rho_{\beta})\leq a$. The prover sends the classical value $\tilde{S}=S(\rho_{\beta})$ and the $E$ register such that the proof state $\chi_{\mathcal{A}E}$ satisfies $\chi_{\mathcal{A}}=\widetilde{I}_{\mathcal{A}}$. [^14] In fact, this is forced as the prover cannot act on the $\mathcal{A}$ register. Furthermore, the prover ensures that the entropy verification protocol on $\chi$ succeeds with probability at least $c$, in which case the average output state $\tilde{\sigma}_{\mathcal{A}}$ satisfies $\|\rho_{\beta}-\tilde{\sigma}\|_{1}\leq\delta$. This is possible because the prover has access to a purification of the $\mathcal{A}$ register and by property 1 of Lemma 3.1.
[…]
For soundness, assume that $\mathcal{F}(H)\geq b$. Let $\chi_{\mathcal{A}E}$ be the proof state and $\tilde{S}$ be the provers entropy claim. If $\chi$ and $\tilde{S}$ are such that the entropy verification protocol accepts with probability less than $s$ there is already a sufficient promise gap with $c-2^{-\Omega{n}}$, the acceptance probability in the YES-case. Suppose thus that the entropy verification protocol accepts with probability at least $s$. Then, by Lemma 3.1 property 2, $S(\tilde{\sigma})\geq\tilde{S}-\delta^{\prime}$. By a Hoeffding bound, the verifiers energy estimate $\tilde{E}$ will satisfy $\tilde{E}\geq\operatorname{Tr}(H\tilde{\sigma})-\delta^{\prime}$ with probability $1-2^{-\Omega(n)}$ (using parallel repetition is necessary). In this case we have:
$\tilde{F}$ $=\tilde{E}-\frac{\tilde{S}}{\beta}$ (76)
$\geq\operatorname{Tr}(H\tilde{\sigma})-\delta^{\prime}-\frac{S(\tilde{\sigma})-\delta^{\prime}}{\beta}$ (77)
$\geq\mathcal{F}-\delta^{\prime}-\frac{\delta^{\prime}}{\beta}$ (78)
$\geq b-\left(1+\frac{1}{\beta}\right)\delta^{\prime}$ (79)
$\geq b-\frac{b-a}{4},$ (80)
which means that the verifier rejects, completing the proof. ∎
6 Low Entropy Ground State Problem
See 1.6
Proof.
We now prove the statement for general Hamiltonians. Treatment of physical Hamiltonians is postponed to Section 8.
We will reduce from the $\mathsf{QMA}(2)$-complete problem $\mathsf{PPIO}$ (Definition 2.7). Let $(U,a,b)$ be a $\mathsf{PPIO}$ instance and assume that $a=2^{-\Omega(n)}$ and $b=2-2^{-\Omega(n)}$. We map this to a $\mathsf{LELES}$ instance $(H_{U},\alpha,\beta,s,t)$ where $H_{U}$ is the Hamiltonian acting on registers $A,B,C$ resulting from the circuit-to-Hamiltonian construction (Lemma 3.2) applied to the channel $\Phi(\rho)\coloneq U(\rho_{A}\otimes\lvert 0^{n_{B}}\rangle\langle 0^{n_{B}}\rvert)U^{\dagger}$. Note that $H_{\Phi}$ is indeed $5$-local. $s$ and $t$ will be $\Omega(n)$ and $O(1)$ respectively and the parameters $\alpha,\beta$ will be fixed at a later point in the proof.
For completeness, let $\left|\psi\right\rangle_{A}$ be such that $\Phi(\psi)$ is $a$-close to $\left|\phi_{A}\right\rangle\otimes\left|\phi_{B}\right\rangle$. By property 1 of Lemma 3.2 there now exists a zero-energy pure state $\left|\chi\right\rangle_{ABC}$ with
$\|\chi_{B}-\lvert\phi_{B}\rangle\langle\phi_{B}\rvert_{B}\|_{1}$ $\leq\|\chi_{B}-\Phi(\psi)_{B}\|_{1}+\|\Phi(\psi)_{B}-\lvert\phi_{B}\rangle\langle\phi_{B}\rvert_{B}\|_{1}$ (81)
$\leq\|\chi_{AB}-\Phi(\psi)_{AB}\|_{1}-\|\Phi(\psi)_{AB}-\lvert\phi_{A}\rangle\langle\phi_{A}\rvert_{A}\otimes\lvert\phi_{B}\rangle\langle\phi_{B}\rvert_{B}\|_{1}$ (82)
$\leq\frac{2T}{T+L+1}+a$ (83)
$\leq O(L^{-1}).$ (84)
Here we used the fact that trace distance is non-increasing under the partial trace. We could absorb $a$ into the $O(L^{-1})$ as $a$ is exponentially small and $L$ will be polynomially large.
Note that $\left|\phi_{B}\right\rangle$ is pure and hence has zero entropy. Using the Fannes Inequality we thus have:
$S(\chi_{B})\leq O\left(\frac{1}{L}(n_{B}+\log L)\right).$ (85)
We now set $\alpha=0$ and $t=O\left(\frac{1}{L}(n_{B}+\log L)\right)$ so that $\left|\chi\right\rangle_{ABC}$ satisfies $\left\langle\chi\right|H_{U}\left|\chi\right\rangle\leq\alpha$ and $S(\chi_{B})\leq t$ as desired. Note that we will later set $L$ sufficiently large such that $t=O(1)$.
For soundness, we first show that any pure state $\left|\psi\right\rangle_{AB}$ that is $b$-far from product has high entanglement entropy. For this, consider the Schmidt decomposition $\left|\psi\right\rangle_{AB}=\sum_{i}\lambda_{i}\left|u_{i}\right\rangle_{A}\left|v_{i}\right\rangle_{B}$. As $\left|\psi\right\rangle_{AB}$ is $b$-far from product, it is in particular $b$-far from $\left|u_{i}\right\rangle\left|v_{i}\right\rangle$ for all $i$ hence we have:
$\frac{1}{2}\|\left|\psi\right\rangle-\left|u_{i}\right\rangle\left|v_{i}\right\rangle\|_{1}$ $=\sqrt{1-|\left\langle u_{i}\right|\left\langle v_{i}\middle|\psi\right\rangle|^{2}}\geq\frac{b}{2}$ (86)
$|\lambda_{i}|^{2}$ $=|\left\langle u_{i}\right|\left\langle v_{i}\middle|\psi\right\rangle|^{2}\leq 1-\frac{b^{2}}{4}$ (87)
Recall the von Neumann entropy is lower bounded by the min-entropy. Hence we have:
$S(\rho_{A})\geq H_{\infty}(\rho_{A})=-\log\|\rho_{A}\|_{\infty}\geq-\log\sqrt{1-\frac{b^{2}}{4}}.$ (88)
Setting b=2xb=2-x yields $S(\rho_{A})\geq-\log\sqrt{x-x^{2}/4}\geq-\frac{1}{2}\log x$. Recall now that $\mathsf{PPIO}$ is $\mathsf{QMA}(2)$-hard even when $b=2-2^{-\Omega(n)}$. We conclude that in the NO-case, all possible isometry outputs must have entropy at least $\Omega(n)$.
Suppose now that for all $\left|\psi\right\rangle$, $\Phi(\psi)$ is $b$ far from any product state. Let $\left|\psi\right\rangle_{ABC}$ be such that $\left\langle\psi\right|H\left|\psi\right\rangle\leq\beta$. By the second property in Lemma 3.2 there exists a pure state $\left|\chi\right\rangle_{A}$ such that
$T\coloneq\|\psi_{AB}-\Phi(\chi)\|_{1}\leq 2\sqrt{\frac{\beta}{\Delta}}+\frac{2T}{T+L+1}.$ (89)
As $\Phi(\chi)$ is far from a product state, we have $S(\Phi(\chi)_{B})\geq\Omega(n)$ by the preceding discussing. By the Fannes Inequality we thus have
$S(\psi_{B})$ $\geq S(\Phi(\chi)_{B})-T(n_{B}-\log T)$ (90)
$\geq\Omega(n)-\left(2\sqrt{\frac{\beta}{\Delta}}+\frac{2T}{T+L+1}\right)\left(n-\log\left(2\sqrt{\frac{\beta}{\Delta}}+\frac{2T}{T+L+1}\right)\right).$ (91)
Setting $\beta=\Delta/L\geq\Omega(L^{-4})$, and choosing $L$ to be a sufficiently large polynomial in $n$ we get $S(\psi_{B})\geq\Omega(n)$ as desired, completing the proof.
7 A $\mathsf{QMA}(2)$-complete local Hamiltonian problem
We will show that depending on the choice of parameters $\alpha,\beta,a,b$, $\mathsf{LEAPS}$ “leaps” from being $\mathsf{QMA}$-complete to being $\mathsf{QMA}(2)$-complete.
See 1.8
Proof.
For containment, note that if $a=0$, then $\mathsf{LEAPS}$ is just a special case of the Separable Local Hamiltonian Problem from [CS12], which is proven to be $\mathsf{QMA}$-complete in the same work. If $a>0$, we can the Hölder Inequality to get
$|\operatorname{Tr}(H\rho)-\operatorname{Tr}(H\sigma)|\leq\|H(\rho-\sigma)\|_{1}\leq\|H\|_{\infty}\|\rho-\sigma\|_{1}.$ (92)
Hence, if $\left\langle\psi\right|H\left|\psi\right\rangle\leq\alpha$ and $\|\psi-\phi_{L}\otimes\phi_{R}\|_{1}\leq a$, then $\left\langle\phi_{R}\right|\left\langle\phi_{L}\right|H\left|\phi_{L}\right\rangle\left|\phi_{R}\right\rangle\leq\alpha+a\|H\|_{\infty}$. To solve the $\mathsf{LEAPS}$ instance $(H,\alpha,\beta,a,b)$ we can now solve the $\mathsf{SeparableLocalHamiltonian}$ instance $(H,\alpha^{\prime},\beta)$ where $\alpha^{\prime}=\alpha+a\|H\|_{\infty}$. Note that $\beta-\alpha^{\prime}\geq\frac{1}{\operatorname{poly}(n)}$ by assumption. Furthermore, note that if $(H,\alpha,\beta,a,b)$ is a YES instance of $\mathsf{LEAPS}$, then $(H,\alpha^{\prime},\beta)$ is a YES instance of $\mathsf{SeparableLocalHamiltonian}$. If $(H,\alpha,\beta,a,b)$ is a NO instance, then certainly all product states have energy $\geq\beta$, hence $(H,\alpha^{\prime},\beta)$ is a NO instance as well. Note that we do not need a gap between $a$ and $b$ as this subsumed by the energy gap.
For hardness we reduce from the Separable Local Hamiltonian problem. Map an instance $(H,\alpha,\beta)$ to the $\mathsf{LEAPS}$ instance $(H,\alpha,\beta^{\prime},0,b)$, where $b\leq\frac{\beta-\alpha}{2\|H\|_{\infty}}$ and $\beta^{\prime}=\beta-\|H\|_{\infty}b$. Note that $\beta^{\prime}-\alpha\geq\frac{\beta-\alpha}{2}\geq\frac{1}{\operatorname{poly}(n)}$. Completeness follows trivially. For soundness, let $\left|\psi\right\rangle$ be a state that is $b$ close to a product state $\left|\phi_{L}\right\rangle\left|\phi_{R}\right\rangle$. By the Hölder inequality, $\left\langle\psi\right|H\left|\psi\right\rangle\geq\left\langle\phi_{R}\right|\left\langle\phi_{L}\right|H\left|\phi_{L}\right\rangle\left|\phi_{R}\right\rangle-\|H\|_{\infty}b\geq\beta-\|H\|_{\infty}b=\beta^{\prime}$, completing the proof. ∎
See 1.9 Note that the regime where we show $\mathsf{QMA}(2)$ completeness has $\beta<a$. As $\|H\|_{\infty}\geq 1$ there is thus not necessarily overlap between the $\mathsf{QMA}$ and $\mathsf{QMA}(2)$ complete regimes. It is in theory possible, although perhaps unlikely, that the $\beta$ and $a$ parameters could be tweaked such that there is overlap between the $\mathsf{QMA}(2)$-hard and contained in $\mathsf{QMA}$ regime. This would prove $\mathsf{QMA}=\mathsf{QMA}(2)$.
Proof.
We first show containment in $\mathsf{QMA}(2)$, then $\mathsf{QMA}(2)$-hardness of general Hamiltonians. The restriction to physical Hamiltonians is postponed to Section 8.
Containment in $\mathsf{QMA}(2)$.
We begin with the characterization $\mathsf{QMA}(k)=\mathsf{QMA}(2)$ from [HM13]. The proofs will be $\left|\psi_{LR}\right\rangle,\left|\phi_{L}\right\rangle$ and $\left|\phi_{R}\right\rangle$. With probability $1/2$ each, the verifier performs one of the following checks:
  1. 1.
    Energy check: the verifier measures the energy [^15] Technically, the energy of $H$ cannot be measured in one go. Here we implicitly use parallel repetition in the standard way. $\left\langle\psi_{LR}\right|H\left|\psi_{LR}\right\rangle$ and accepts if this is $\frac{\alpha+\beta}{2}$.
  2. 2.
    Distance check: the verifier uses the SWAP test between $\left|\psi_{LR}\right\rangle$ and $\left|\phi_{L}\right\rangle\left|\phi_{R}\right\rangle$ and accepts if the SWAP test accepts.
To see completeness, the provers send $\left|\psi_{LR}\right\rangle,\left|\phi_{L}\right\rangle,\left|\phi_{R}\right\rangle$ such that $\left\langle\psi_{LR}\right|H\left|\psi_{LR}\right\rangle\leq\alpha$ and $\|\left|\psi_{LR}\right\rangle-\left|\phi_{L}\right\rangle\left|\phi_{R}\right\rangle\|_{1}\leq a$. The energy check can be made to succeed with probability $1-2^{-\Omega(n)}$ using parallel repetition. The distance check will accept with probability
$p_{acc}=\frac{1}{2}+\frac{1}{2}\left|\left\langle\psi_{LR}\middle|\phi_{L}\right\rangle\left|\phi_{R}\right\rangle\right|^{2}=1-\frac{1}{8}\|\psi_{LR}-\phi_{L}\otimes\phi_{R}\|_{1}^{2}\geq 1-\frac{a^{2}}{8}.$ (93)
The total acceptance probability is therefore at least $1-2^{-\Omega(n)}-\frac{a^{2}}{8}\geq\frac{1}{2}+\frac{1}{\operatorname{poly}(n)}$, where we use the fact that $a\leq b-\frac{1}{\operatorname{poly}(n)}\leq 2-\frac{1}{\operatorname{poly}(n)}$.
For soundness, note that if $\left\langle\psi_{LR}\right|H\left|\psi_{LR}\right\rangle\geq\beta$, then the energy check will fail with probability inverse exponentially close to 1. Hence, if the provers send such $\left|\psi_{LR}\right\rangle$, the verifier accepts with probability at most $\frac{1}{2}+2^{-\Omega(n)}$, which is sufficiently far from the acceptance probability in the YES case to distinguish both cases. We conclude we can assume $\left\langle\psi_{LR}\right|H\left|\psi_{LR}\right\rangle<\beta$, from which it follows that $\|\psi_{LR}-\phi_{L}\phi_{R}\|_{1}>b$ as we are in the NO case. The acceptance probability of the distance check will now be at most $1-\frac{b^{2}}{8}$, which is inverse polynomially bounded away from $1-2^{-\Omega(n)}-\frac{a^{2}}{8}$, the acceptance probability in the YES case. This completes the soundness proof.
$\mathsf{QMA}(2)$-hardness.
We will reduce from the $\mathsf{QMA}(2)$-complete problem $\mathsf{PPIO}$ (Definition 2.7). Let $(U,a^{\prime},b^{\prime})$ be a $\mathsf{PPIO}$ instance where $U$ is a unitary acting on registers $A,B$ of $n_{A}+n_{B}=n$ qubits. Assume that $a^{\prime}=2^{-\Omega(n)}$ and $b^{\prime}=2-2^{-\Omega(n)}$. We again use the channel-to-Hamiltonian construction from Lemma 3.2, this time applied to the channel $\Phi(\rho)\coloneq\operatorname{Tr}_{A}\left(U(\rho_{A}\otimes\lvert 0^{n_{B}}\rangle\langle 0^{n_{B}}\rvert)U^{\dagger}\right)$. Let $H_{U}$ be the resulting Hamiltonian acting on registers $A,B,C$. The output of the reduction will be $(H_{U},\alpha,\beta,a,b)$ with $\alpha=0,\beta=\Omega(L^{-3}),a=O(L^{-1/2})$ and $b=O(1)$. Note that $H_{U}$ is indeed $5$-local.
For completeness, let $\left|\psi\right\rangle,\left|\phi_{L}\right\rangle,\left|\phi_{R}\right\rangle$ be such that $\|U(\psi\otimes\lvert 0^{n_{B}}\rangle\langle 0^{n_{B}}\rvert)U^{\dagger}-\phi_{L}\otimes\phi_{R}\|\leq a$. Trace distance is non-increasing under partial trace, so in particular $\|\Phi(\psi)-\phi_{R}\|_{1}\leq a$. By Lemma 3.2 property 1 there now exists a zero-energy pure state $\left|\chi\right\rangle$ such that $\|\chi_{B}-\Phi(\psi)\|_{1}\leq\frac{2T}{T+L+1}+a$. By the Gentle Measurement Lemma, it follows that $\chi_{ABE}$ is close to a product state. More precisely, there is some $|\tilde{\phi}\rangle_{AC}$ such that
$\|\chi_{ABC}-\tilde{\phi}_{AC}\otimes\phi_{R,B}\|$ $\leq 2\sqrt{\frac{2T}{T+L+1}+a}$ (94)
$\leq O\left(\frac{1}{\sqrt{L}}\right),$ (95)
where in the last step we use that $L$ is polynomial in $n$ to absorb the $a^{\prime}=2^{-\Omega(n)}$. Setting $a=\Omega\left(\frac{1}{\sqrt{L}}\right)$ we see that $\left|\chi\right\rangle$ has zero energy and is $a^{\prime}$-close to a product state as desired.
For soundness, assume $(U,a^{\prime},b^{\prime})$ is a NO instance and let $\left|\psi\right\rangle_{ABC}$ be an arbitrary state with $\left\langle\psi\right|H\left|\psi\right\rangle\leq\beta$. By Lemma 3.2 property 2, there exists a pure state $\left|\chi\right\rangle_{A}$ such that $\|\psi_{B}-\Phi(\chi)\|_{1}\leq\sqrt{\beta/\Delta}+\frac{2T}{T+L+1}$. Suppose now, towards a contradiction, that $\left|\psi\right\rangle$ is close to product: suppose there are $\left|\phi_{AC}\right\rangle,\left|\phi_{B}\right\rangle$ such that $\|\psi_{ABC}-\phi_{AC}\otimes\phi_{B}\|_{1}\leq b$. By monotonicity under partial trace of trace distance we have:
$\|\Phi(\chi)-\phi_{B}\|_{1}$ $\leq\|\Phi(\chi)-\psi_{B}\|_{1}+\|\psi_{B}-\phi_{B}\|_{1}$ (96)
$\leq\|\Phi(\chi)-\psi_{B}\|_{1}+\|\psi_{ABC}-\phi_{AC}\otimes\phi_{B}\|_{1}$ (97)
$\leq\sqrt{\beta/\Delta}+\frac{2T}{T+L+1}+b.$ (98)
As we are in the NO case, $U\left|\psi\right\rangle\otimes\left|0^{n_{B}}\right\rangle$ is $b^{\prime}$ far from any pure state. In particular, by monotonicity of trace distance, $\|\Phi(\chi)-\phi_{B}\|_{1}\geq b^{\prime}$. We have thus achieved a contradiction if
$\sqrt{\beta/\Delta}+\frac{2T}{T+L+1}+b<2-2^{-\Omega(n)}\leq b^{\prime}.$ (99)
In particular, setting $\beta\leq\Delta$, $b=O(1)$ and $L=n^{\Omega(1)}$ achieves $\sqrt{\beta/\Delta}+\frac{2T}{T+L+1}+b\leq O(1)$ and hence suffices. Note that $\Delta=\Omega(L^{-3})$, hence setting $\beta=O(L^{-3})$ suffices. We now have $a=\Omega\left(\frac{1}{\sqrt{L}}\right)$ and $\beta=O\left(\frac{1}{L^{3}}\right)$, hence $\beta=O(a^{6})$. ∎
8 Physical Hamiltonians
In this section we will use the universal Hamiltonian simulator framework developed in [CMP18, ZA21] to extend our results to physically motivated Hamiltonians. The main idea of this framework is to encode any Hamiltonian $H$ into the low energy space of a bigger Hamiltonian $H^{\prime}$, which belongs to some set of physically relevant Hamiltonians of choice, such as Heisenberg Hamiltonians or $XY$-interaction Hamiltonians. We use a simplified notion of Hamiltonian simulation than that originally introduced in [CMP18].
Definition 8.1 (Hamiltonian simulation (simplified version of [CMP18, Definition 23])).
An $n$-qudit Hamiltonian $H$ is $(\Delta,\eta,\varepsilon)$-simulated by $H^{\prime}$ if for some local isometry $V=\bigotimes_{i}V_{i}$, where each $V_{i}$ maps a single qubit to $O(1)$ qubits, there is a (not necessarily local) isometry $\widetilde{V}$ such that
  1. 1.
    $V$ is close to $\widetilde{V}$: $\|\widetilde{V}-V\|_{\infty}\leq\eta$.
  2. 2.
    $\widetilde{V}$ maps the identity to the projector on the subspace of states with energy $\leq\Delta$ according to $H^{\prime}$. That is, $\widetilde{V}I\widetilde{V}^{\dagger}=P_{\leq\Delta(H^{\prime})}$
  3. 3.
    The low energy part of the spectrum of $H^{\prime}$ is close to the spectrum of $H$. That is, $\|H^{\prime}_{\leq\Delta}-\widetilde{V}H\widetilde{V}^{\dagger}\|_{\infty}\leq\varepsilon$, where $H^{\prime}_{\leq\Delta}=H^{\prime}P_{\leq\Delta(H^{\prime})}$.
We say the simulation is efficient if the description of $H^{\prime}$ can be efficiently computed and both number of particles $H^{\prime}$ acts on, and its maximum energy $\|H^{\prime}\|_{\infty}$ are polynomial in $n,\eta^{-1},\varepsilon^{-1},\Delta$.
Definition 8.2 (Universal family of Hamiltonians ([CMP18, ZA21])).
A family $\mathcal{F}=\{H_{i}\}$ is weakly universal if for any $\Delta,\eta,\varepsilon>0$, any $O(1)$-local Hamiltonian $H$ can be $(\Delta,\eta,\varepsilon)$-simulated by some $H_{i}\in\mathcal{F}$. We say the family is strongly universal if this simulation is always efficient.
The main result of [ZA21] says that most families of Hamiltonians on the 2D square lattice are strongly universal. The exception are families that are 2SLD, meaning that their 2-local interactions can be simultaneously and locally diagonalized. More precisely:
Definition 8.3 (2SLD ([CMP18])).
Let $\mathcal{S}$ be a set of 2-local interaction terms on qubits. $\mathcal{S}$ is 2SLD if there exists $U\in\mathcal{SU(2)}$ such that $U^{\otimes 2}H_{i}(U^{\dagger})^{\otimes 2}=\alpha_{i}Z\otimes Z+A_{i}\otimes I+I\otimes B_{i}$ for all $H_{i}\in\mathcal{S}$. Here $\alpha_{i}\in\mathbb{R}$ and $A_{i},B_{i}$ are arbitrary 1-local operators.
Originally, [ZA21] uses the fully general defintion of Hamiltonian simulation. However, closer inspection of their results reveals that their simulations have the form of Definition 8.1.
Theorem 8.4 (Main theorem of [ZA21]).
Let $\mathcal{S}$ be a set of 2-qubit interactions and $\mathcal{F}(\mathcal{S})$ be the family of Hamiltonians on a 2D square lattice where all 2-qubit interactions are from $\mathcal{S}$, with arbitrary interaction strenghts. If $\mathcal{S}$ is non-2SLD, then $\mathcal{F}(\mathcal{S})$ is strongly universal.
Corollary 8.5.
The families of Heisenberg Hamiltonians ($\mathcal{S}=\{X\otimes X+Y\otimes Y+Z\otimes Z\}$) and $XY$-interaction Hamiltonians ($\mathcal{S}=\{X\otimes X+Y\otimes Y\}$) are strongly universal.
Using this framework we easily get that all hardness results in this work hold even when restricted to any non-2SLD family of Hamiltonians.
Corollary 8.6.
Let $\mathcal{S}$ be a non-2SLD set of $2$-qubit interactions and let $\mathcal{F}(\mathcal{S})$ be as in Theorem 8.4. Then
  1. 1.
    $\mathsf{HELES}$ is $\mathsf{qq\mhyphen QAM}$-hard, even when restricted to $\mathcal{F}(\mathcal{S})$.
  2. 2.
    $\mathsf{LELES}$ is $\mathsf{QMA}(2)$-hard, even when restricted to $\mathcal{F}(\mathcal{S})$
  3. 3.
    $\mathsf{LEAPS}$ is $\mathsf{QMA}(2)$-hard, even when restricted to $\mathcal{F}(\mathcal{S})$
Proof.
We will first consider $\mathsf{HELES}$. The proof for $\mathsf{LELES}$ will be completely analogous.
Let $(H_{AB},\alpha,\beta,s,t)$ be a $\mathsf{HELES}$ instance and let $H^{\prime}_{A^{\prime}B^{\prime}}$ $(\Delta,\eta,\varepsilon)$-simulate $H$ via the local isometry $V$. We map our instance to $(H^{\prime}_{A^{\prime}B^{\prime}},\alpha^{\prime},\beta^{\prime},s^{\prime},t^{\prime})$ where $\alpha^{\prime}=\alpha+\frac{\beta-\alpha}{4},\beta^{\prime}=\beta-\frac{\beta-\alpha}{4},s^{\prime}=s$ and $t^{\prime}=t+\frac{s-t}{4}$. We now show that YES and NO instances are preserved.
For completeness, let $\left|\psi\right\rangle_{AB}$ be a state with $\left\langle\psi\right|H\left|\psi\right\rangle\leq\alpha$. Consider the state $\left|\psi^{\prime}\right\rangle=V\left|\psi\right\rangle$. Note that $\|H^{\prime}-VHV^{\dagger}\|_{\infty}\leq\|H\|_{\infty}\|\widetilde{V}-V\|_{\infty}\leq\eta\|H\|_{\infty}$. Therefore, $\left\langle\psi^{\prime}\right|H^{\prime}\left|\psi^{\prime}\right\rangle\leq\alpha+\|H\|_{\infty}\eta$. Choosing $\eta\leq\frac{\beta-\alpha}{4\|H\|_{\infty}}$ we get that $\left\langle\psi^{\prime}\right|H^{\prime}\left|\psi^{\prime}\right\rangle\leq\alpha^{\prime}$ is low energy. Furthermore, as $V$ is a local isometry it preserves entropy. In particular, $S(\psi_{A})=S(\psi^{\prime}_{A^{\prime}})$.
For soundness suppose $\left|\psi^{\prime}\right\rangle$ satisfies $\left\langle\psi^{\prime}\right|H^{\prime}\left|\psi^{\prime}\right\rangle\leq\beta^{\prime}$. Let $\left|\phi^{\prime}\right\rangle$ be the projection of $\left|\psi^{\prime}\right\rangle$ on the low energy space of $H^{\prime}$. Note that
$\beta$ $\geq\left\langle\psi^{\prime}\right|H^{\prime}\left|\psi^{\prime}\right\rangle$ (100)
$\geq\Delta\left\langle\psi^{\prime}\right|(I-\Pi_{\leq\Delta})\left|\psi^{\prime}\right\rangle$ (101)
$\geq\Delta(1-\left\langle\psi^{\prime}\right|\Pi_{\leq\Delta}\left|\psi^{\prime}\right\rangle)$ (102)
$\left\langle\psi^{\prime}\right|\Pi_{\leq\Delta}\left|\psi^{\prime}\right\rangle$ $\geq 1-\frac{\beta}{\Delta},$ (103)
hence by the Gentle Measurement Lemma, $\|\psi^{\prime}-\phi^{\prime}\|_{1}\leq 2\sqrt{\frac{\beta}{\Delta}}$.
[…]
$\left\langle\phi\right|H\left|\phi\right\rangle$ $\leq\left\langle\phi^{\prime}\right|\widetilde{V}H\widetilde{V}^{\dagger}\left|\phi^{\prime}\right\rangle$ (104)
$\leq\left\langle\phi^{\prime}\right|H^{\prime}\left|\phi^{\prime}\right\rangle+\varepsilon$ (105)
$\leq\left\langle\psi^{\prime}\right|H^{\prime}\left|\psi^{\prime}\right\rangle+2\|H\|_{\infty}\sqrt{\frac{\beta}{\Delta}}+\varepsilon.$ (106)
Hence, for an appropriate choices of $\Delta,\varepsilon$, $\left\langle\phi\right|H\left|\phi\right\rangle\leq\beta$. As we are in the NO case, this means that $S(\phi_{A})\leq t$ and as $V$ is a local isometry, $S(\chi_{A})\leq t$ for $\left|\chi\right\rangle=V\left|\phi\right\rangle$. Note that $\|\chi-\phi^{\prime}\|_{1}\leq\|V-\tilde{V}\|_{\infty}=\eta$ and that $\|\phi^{\prime}-\psi^{\prime}\|_{1}\leq 2\sqrt{\frac{\beta}{\Delta}}$, hence $\|\chi-\psi^{\prime}\|_{1}\leq\eta+2\sqrt{\frac{\beta}{\Delta}}$. By the Fannes Inequality, the entropy of $\psi^{\prime}_{A^{\prime}}$ is now close to that of $\phi_{A}$. In particular, for appropriate choices of $\eta,\Delta$, we get $S(\psi^{\prime}_{A^{\prime}})\leq t^{\prime}$.
The proof for $\mathsf{LELES}$ follows by swapping the roles of $t$ and $s$ and the corresponding inequalities.
For $\mathsf{LEAPS}$, the situation is a bit more subtile as, at least in principle, there could be product states that are only product because of a very small contribution of the high energy space of $H^{\prime}$. Nevertheless, the proof still proceeds largely along the same lines as $\mathsf{HELES}$. We map our $\mathsf{LEAPS}$ instance $(H,\alpha,\beta,a,b)$ to $(H^{\prime},\alpha^{\prime},\beta^{\prime},a^{\prime},b^{\prime})$ where $H^{\prime},\alpha^{\prime},\beta^{\prime}$ are as above. For the $a,b$ parameters we will set $a^{\prime}=a=O(n^{-3})$, $b=\Omega(1)$ and $b^{\prime}=\Theta(n^{-2})$. Note that the $\mathsf{QMA}(2)$-hardness proof of $\mathsf{LEAPS}$ indeed allows these choices of $a,b$.
For completeness it sufficies to note that $V\left|\phi\right\rangle_{AB}$ is at most as from product as $\left|\phi\right\rangle_{AB}$ since $V$ is a local isometry.
For completeness, we need to show that if $\left|\phi\right\rangle_{AB}$ is far from product, then so is $\left|\psi^{\prime}\right\rangle_{A^{\prime}B^{\prime}}$. Towards a contradiction, assume there exist $\left|\theta\right\rangle_{A^{\prime}},\left|\theta\right\rangle_{B^{\prime}}$ such that $\|\psi^{\prime}_{A^{\prime}B^{\prime}}-\theta_{A^{\prime}}\otimes\theta_{B^{\prime}}\|_{1}\leq b^{\prime}$. It follows that $\|\phi^{\prime}_{A^{\prime}B^{\prime}}-\theta_{A^{\prime}}\otimes\theta_{B^{\prime}}\|_{1}\leq b^{\prime}+2\sqrt{\frac{\beta}{\Delta}}\eqcolon d$. By the Fannes Inequality, we now have $S(\phi^{\prime}_{A^{\prime}})\leq nd-d\log d$. As $V$ is a local isometry it preserves entropy, hence $S(\phi_{A})\leq nd-d\log d$ as well.
Just like before we have $\left\langle\phi\right|H\left|\phi\right\rangle\leq\beta$ by our choice of parameters. As we are in the NO case, this implies that $\left|\phi\right\rangle_{AB}$ is $b$-far from any product state. In particular, $\left|\phi\right\rangle_{AB}$ is $b$-far from the product states in its own Schmidt decomposition, and $S(\phi_{A})\geq-\log\sqrt{1-\frac{b^{2}}{4}}$ (similary to Eq. 88). As we have $b=\Omega(1)$, we get our desired contradition if $nd-d\log d=o(1)$. The proof now follows by setting $\Delta=\left(\frac{\beta}{n^{2}}\right)$ so that $d=O(n^{-2})$.
Acknowledgements.
The authors would like to thank Dorian Rudolph for bringing the class $\mathsf{qq\mhyphen QAM}$ to our attention.